### v1

# Imitation Learning for Combinatorial Optimisation under Uncertainty

Abstract
Imitation learning (IL) provides a data-driven framework for approximating policies for large-scale combinatorial optimisation problems formulated as sequential decision problems (SDPs), where exact solution methods are computationally intractable. A central but underexplored aspect of IL in this context is the role of the expert that generates training demonstrations. Existing studies employ a wide range of expert constructions, yet lack a unifying framework to characterise their modelling assumptions, computational properties, and impact on learning performance.
This paper introduces a systematic taxonomy of experts for IL in combinatorial optimisation under uncertainty. Experts are classified along three dimensions: (i) their treatment of uncertainty, including myopic, deterministic, full-information, two-stage stochastic, and multi-stage stochastic formulations; (ii) their level of optimality, distinguishing task-optimal and approximate experts; and (iii) their interaction mode with the learner, ranging from one-shot supervision to iterative, interactive schemes. Building on this taxonomy, we propose a generalised Dataset Aggregation (DAgger) algorithm that supports multiple expert queries, expert aggregation, and flexible interaction strategies.
The proposed framework is evaluated on a dynamic physician-to-patient assignment problem with stochastic arrivals and capacity constraints. Computational experiments compare learning outcomes across expert types and interaction regimes. The results show that policies learned from stochastic experts consistently outperform those learned from deterministic or full-information experts, while interactive learning improves solution quality using fewer expert demonstrations. Aggregated deterministic experts provide an effective alternative when stochastic optimisation becomes computationally challenging.
keywords:
Imitation Learning, Operations Research, Sequential Decision Making
category:
[…]
1 Introduction
Combinatorial optimisation (CO) lies at the core of many real-world decision-making problems, including scheduling, logistics, network design, and resource allocation. These problems are challenging due to the combinatorial explosion in the number of feasible solutions, whereby even relatively small instances can generate an enormous search space. This difficulty is further compounded by complex constraints and the requirement to identify globally optimal solutions rather than merely feasible ones. As problem size increases, naïve or brute-force approaches quickly become impractical, necessitating more sophisticated methods for efficiently exploring the solution space.
The computational burden becomes particularly severe in settings involving uncertainty or temporal dynamics, such as stochastic or sequential decision problems, where multiple future scenarios must be considered. In such cases, exact optimisation methods (such as branch-and-bound or dynamic programming) often become computationally intractable. This challenge is further exacerbated in applications that require fast and reliable decision-making. To address these limitations, practitioners rely on a variety of approximation techniques, including heuristics, metaheuristics, and problem-specific algorithms. These approaches trade off optimality guarantees for computational efficiency, thereby enabling the practical deployment of CO methods in complex operational environments.
Machine learning (ML) has recently emerged as a powerful complementary tool for tackling CO problems, particularly when traditional techniques scale poorly. In high-dimensional or uncertain environments, ML methods can learn approximate decision rules by exploiting patterns in data, enabling faster and more adaptive decision making. This capability is especially valuable in real-time operational settings. Reinforcement learning (RL) [sutton1998reinforcement], in particular, has demonstrated strong potential by learning decision policies through interaction with an environment, using reward feedback to iteratively improve performance. Such approaches are well-suited to settings where explicit models are unavailable, incomplete, or prohibitively complex.
Alongside RL, Imitation learning (IL) [hussein2017imitation] has gained attention as an attractive alternative, especially in scenarios where defining a suitable reward function is difficult but expert demonstrations are available. Instead of learning through trial and error, IL aims to learn a direct mapping from system states to actions by mimicking expert behaviour. A central challenge in this paradigm is the acquisition of high-quality expert demonstrations, which are often expensive or infeasible to generate due to the computational cost of solving CO problems optimally. This has motivated growing interest in approximate, surrogate, or interactive expert models that can provide informative guidance while remaining computationally tractable. The design and utilisation of such experts remain key open questions at the intersection of ML and combinatorial optimisation.
Contribution
The concept of imitating an expert policy within CO is relatively recent, and existing work proposes a wide range of approaches for defining or approximating experts. These include offline optimal solutions, simplified surrogate models, and structured or repeated interactions with problem-specific oracles. However, to the best of our knowledge, the literature lacks a systematic understanding of what constitutes an expert and how expert choices influence learning outcomes in IL for CO. To address this gap, this paper introduces a taxonomy of expert design for IL in CO. The taxonomy categorises experts along key dimensions, providing a conceptual framework to guide the principled selection and design of experts in different problem settings.
To support this taxonomy, we present a survey of recent literature that illustrates the diversity of expert construction and usage strategies. In the second part of the paper, we propose a two-stage learning pipeline in which a CO-based expert layer generates demonstrations that are subsequently used to train an ML-based decision model. By systematically varying expert properties within this framework, we conduct a comparative analysis that highlights the relative strengths and limitations of different expert types in the context of a physician scheduling problem. The results provide practical insights into how expert design choices affect learning performance, computational cost, and solution quality in sequential decision-making problems.
Scope
Several existing surveys examine the integration of ML and CO [bengio2021machine, kotary2021end, mandi2024decision, sadana2025survey, scavuzzo2024machine]. In contrast, this article focuses specifically on the role and design of the expert in IL. We concentrate on settings involving sequential decision making under uncertainty, where expert queries are costly and real-time decisions are required. Nevertheless, we also consider studies involving static and two-stage stochastic problems for completeness. Experts may take the form of human decision makers, problem-specific heuristics, or exact optimisation solvers. The overarching objective is to develop learning-based models that approximate or replace expert decision making, thereby enabling scalable and timely decisions without substantial loss in solution quality. No assumptions are made regarding specific structural properties of the underlying optimisation problems.
Organisation
The remainder of the paper is organised as follows. Section 2 reviews key concepts in ML and CO for sequential decision making. Section 3 introduces the fundamentals of IL. Section 4 presents the proposed taxonomy of expert policies and situates it within the existing literature. Section 5 describes the problem setting considered in this study, and Section 6 details the proposed methodology for applying IL to sequential decision making under uncertainty. Section 7 presents the experimental results and analysis. Section 8 concludes the paper.
2 Background
Many real-world combinatorial optimisation problems can be naturally formulated as sequential decision problems (SDPs) [powell2007approximate]. At each decision epoch, an agent selects an action based on the current system state and the information available at that time, while accounting for the impact of the decision on both immediate outcomes and future system evolution. The decision process is inherently dynamic, as new information becomes available over time and the system evolves in response to previously taken actions. The objective is typically to design policies that minimise cumulative cost or maximise cumulative reward over a finite or infinite planning horizon. This modelling framework is particularly well suited to applications that involve uncertainty, temporal structure, and adaptive decision-making.
Sequential decision problems are commonly modelled using the Markov decision process (MDP) framework. We begin by introducing this framework in general terms, outlining its key components and underlying assumptions. We then briefly review exact solution methods, including dynamic programming, and discuss their primary limitation, namely, computational intractability in large-scale problems due to the curse of dimensionality. To support the developments in later sections, we also introduce fundamental concepts from machine learning that are relevant to sequential decision making. Finally, we discuss how machine learning techniques have been employed to develop approximate solution methods for MDPs, offering scalable alternatives when exact approaches are computationally infeasible.
2.1 Markov Decision Process
A Markov decision process is a mathematical framework used to model decision-making problems in which outcomes are influenced by both random factors and controllable actions [powell2022reinforcement]. The core elements of an MDP are as follows.
  • Decision epoch ($k$) - A decision epoch is a point in time at which the decision maker observes the current system state and selects an action. Let $K$ denote the final decision epoch.
  • State ($x_{k}$) - The state provides a complete description of the system at decision epoch $k$ that is sufficient for making future decisions. The collection of all possible states defines the state space $\mathcal{S}$. When the epoch index is not required, we use $x$ to denote a generic state.
  • Decision variables ($a_{k}$) - A decision variable represents the action taken at epoch $k$. It may be binary, discrete, continuous, or vector-valued, and is subject to feasibility constraints of the form $a_{k}\in A_{k}$.
  • Exogenous information ($\xi_{k}$) - Exogenous information represents uncertainty revealed after a decision is made. This information is not known at the time of decision-making and originates from outside the system being modelled.
  • Cost/Reward Function ($C_{k}(x_{k},a_{k})$) – This function specifies the cost incurred or reward obtained when action $a_{k}$ is taken in state $x_{k}$ at decision epoch $k$.
  • Transition function ($X^{M}(x_{k},a_{k},\xi_{k})$) - The transition function specifies how the system evolves from one state to the next. It defines the equations used to update each component of the state and captures all system dynamics, including the updating of estimates and beliefs.
  • Policy ($\pi$) - A policy is a mapping from a state $x_{k}$ to an admissible action $a_{k}$. Policies may be represented in tabular form or as parameterised functions, for example, using parameters $\theta$. The optimal policy $\pi^{*}$ is defined as
    $\pi^{*}$ $=\operatorname*{arg\,min}_{\pi}\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},\pi(x_{k}))\,\big|\,x_{0}\Bigg\}$ (1)
    $=\operatorname*{arg\,min}_{\pi}V^{\pi}(x_{k}).$ (2)
  • Value function - The value function represents the cumulative cost or reward accrued over the planning horizon. Assuming a cost-minimisation objective, the value function from an initial state $x_{0}$ is given by
    $V^{\pi^{*}}(x_{0})=\min_{\pi}\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},\pi(x_{k}))\,\big|\,x_{0}\Bigg\},$ (3)
    where $a_{k}=\pi(x_{k})$. For simplicity, discounting of future costs is omitted, with a primary focus on finite-horizon problems. In the MDP literature, $V^{\pi^{*}}(x_{k})$ is referred to as the optimal value function or optimal cost-to-go from state $x_{k}$ at epoch $k$.
    Given a value function, the optimal policy can be recovered via a one-step greedy minimisation. Value functions for a fixed policy can be computed using dynamic programming methods [puterman2014markov], and Equation˜2 can be solved using value iteration or policy iteration algorithms. Value iteration repeatedly updates state values using the Bellman optimality equation until convergence, after which the optimal policy is derived. Policy iteration alternates between policy evaluation and policy improvement steps. A third approach formulates the MDP as a linear programme, optimising the value function subject to Bellman constraints. While these methods yield exact solutions, they are computationally expensive and often infeasible for large-scale problems due to the three curses of dimensionality [powell2007approximate].
    In some applications, interest lies only in the optimal action at a specific epoch $k$. The optimal action $a_{k}^{*}$ is given by
    $a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\Big\{C_{k}(x_{k},a_{k})+\mathbb{E}[V_{k+1}(x_{k+1})]\,\big|\,x_{k}\Big\},$ (4)
    where $V_{k+1}(x_{k+1})$ denotes the optimal value of the next state.
The nature of a decision-making problem depends on the length of the planning horizon $K$. When $K<\infty$, the problem is finite-horizon, whereas $K=\infty$ corresponds to an infinite-horizon setting. A particularly important special case arises when $K=1$. When uncertainty is present, this setting corresponds to a two-stage stochastic programming problem with a single realisation of uncertainty. In the absence of uncertainty, it reduces to a deterministic optimisation problem. Although such problems are static in nature, they often arise repeatedly in practice. For example, a company may solve an independent delivery routing problem each day, with no direct dependence across days. When $K>1$, the problem becomes a multi-stage or fully sequential decision problem.
2.2 Machine Learning
Machine learning (ML) is a subfield of artificial intelligence that focuses on designing algorithms capable of learning patterns from data to make predictions or decisions without explicit programming. At a high level, the objective is to infer a function that maps inputs to corresponding outputs. Formally, let $X$ denote the input space, commonly referred to as the feature space, and let $Y$ denote the output space. Each input encodes a collection of descriptive attributes for a data instance, and the task is to produce an appropriate output based on these features. We consider a class of parameterised models represented by functions $f_{\theta}:X\rightarrow Y$, where $\theta\in\Theta$ denotes the set of model parameters. The learning process consists of selecting parameter values that yield an effective input–output mapping, typically guided by empirical data and a chosen performance criterion. Depending on the type of feedback available, learning problems are commonly categorised into three main settings: supervised learning, where labelled data are provided; unsupervised learning, where only unlabelled data are available; and reinforcement learning, where an agent learns through interaction with an environment and receives feedback in the form of rewards.
Supervised learning addresses settings in which the learner is provided with a finite collection of input–output pairs $\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}\in X$ and $y_{i}\in Y$ denote the input and corresponding output, respectively. The outputs $y_{i}$ are commonly referred to as labels or ground truth. The objective is to approximate the underlying input–output relationship by minimising a loss function $\ell:Y\times Y\rightarrow\mathbb{R}$, which measures the discrepancy between predicted and observed outputs. A widely used example is the mean squared error (MSE), defined as
$\frac{1}{N}\sum_{i=1}^{N}\left(f_{\theta}(x_{i})-y_{i}\right)^{2}.$
Since the true data-generating distribution is unknown, the loss is minimised empirically over the observed samples.
Unsupervised learning, by contrast, refers to settings in which only input samples $\{x_{i}\}_{i=1}^{N}$ are available, without associated labels. The goal is to uncover latent structure or meaningful representations within the data. Typical tasks include clustering, dimensionality reduction, and generative modelling. These methods rely on objective functions that promote internal consistency or compactness of the learned representations, such as minimising intra-cluster distances in $k$-means clustering. In the absence of ground-truth labels, evaluation typically relies on internal metrics or qualitative assessments informed by domain knowledge.
Reinforcement learning concerns sequential decision-making problems in which a learning agent interacts with an environment over time. At each time step, the agent observes a state, selects an action according to a policy, and receives feedback in the form of a scalar reward, along with a new state resulting from its action. Unlike supervised learning, explicit target outputs are not provided; instead, the agent must infer desirable behaviour from delayed and often sparse rewards. The objective of reinforcement learning is to learn a policy that maximises the expected cumulative reward over time, typically formalised using frameworks such as MDPs. This setting is particularly well-suited to problems involving control, planning, and adaptation, where actions influence future states and feedback is evaluative rather than instructive.
The learning process involves several interconnected components. First, a model class is selected, such as linear predictors, decision trees, kernel methods, or neural networks, each offering different expressive capabilities. Learning proceeds by specifying an appropriate loss function and employing an optimisation algorithm, such as gradient descent or its variants, to adjust model parameters to minimise the loss on the training data. In addition to model parameters, learning algorithms involve hyperparameters, such as learning rates, regularisation strengths, or architectural choices, which govern the learning dynamics and model complexity. To evaluate generalisation performance and mitigate overfitting, data are typically partitioned into a training set, a validation set for hyperparameter tuning, and a test set for final evaluation. Overfitting occurs when a model fits the training data well but fails to generalise to unseen instances.
2.3 Machine Learning for MDPs
Machine learning provides a range of approaches for solving MDPs in settings where system dynamics are partially known or entirely unknown. RL and approximate dynamic programming (ADP) are two prominent frameworks in this context. RL is commonly applied in model-free settings where transition dynamics are unknown. Algorithms such as Q-learning and policy gradient methods learn policies directly through interaction with the environment, using sampled trajectories to estimate value functions or policy gradients and to iteratively improve performance.
Approximate dynamic programming, more prevalent in operations research and control, often assumes access to a simulator or generative model of the system. ADP addresses the intractability of exact dynamic programming by approximating value functions or policies using function approximation techniques, including linear basis expansions, regression trees, and neural networks. These approximations enable iterative solution methods such as fitted value iteration, approximate policy iteration, and least-squares temporal difference learning. Although RL and ADP differ in historical origins and terminology, the distinction between the two has become increasingly blurred, particularly with the adoption of deep learning–based methods.
Imitation learning (IL) offers an alternative approach for solving MDPs, especially in situations where the reward function is unknown, difficult to specify, or poorly aligned with the desired behaviour. In this setting, the objective is to learn a policy by leveraging expert-generated state–action trajectories. The simplest form of IL is Behavioural Cloning (BC) [pomerleau1988alvinn], which treats imitation as a supervised learning problem that maps observed states directly to expert actions. However, BC is vulnerable to compounding errors arising from distributional shift, whereby the learner encounters states during deployment that are underrepresented in the training data.
To address this issue, more robust IL methods incorporate expert interaction during learning, allowing the learner to query the expert on states encountered during execution. A related class of methods, known as inverse reinforcement learning (IRL) [ng2000algorithms], seeks to infer a reward function for which the expert policy is optimal. Once such a reward function is recovered, standard reinforcement learning or approximate dynamic programming techniques may be applied to compute a policy.
The focus of this work is the application of IL to CO problems under uncertainty, with particular emphasis on the role and design of the expert. Building on the formulation of MDPs and the use of expert demonstrations, the following section provides a detailed overview of IL techniques.
3 Imitation Learning: Models and Challenges
Imitation learning (IL) [hussein2017imitation] is a specialised branch of supervised machine learning in which an agent learns a task by mimicking behaviour demonstrated by an expert, to replicate the expert decision-making process from observed examples. Unlike classical supervised learning, where data points are typically assumed to be independent and identically distributed, IL often involves sequential decision making. In such settings, data points are temporally correlated and depend on previously taken actions.
The term expert denotes the source of demonstrations used by the learning agent and may correspond to a human decision maker or an exact algorithm. Demonstrations are commonly provided as trajectories $\tau=\{x_{i},a_{i}^{*}\}_{i=1}^{T}$, where $T$ denotes the trajectory length. An expert dataset can then be defined as $\mathcal{D}=\{\tau_{j}\}_{j=1}^{N}$, where each state $x_{i}$ is generated under the expert policy, that is $x_{i}\sim\pi^{*}$, and the corresponding action is given by $a_{i}^{*}=\pi^{*}(x_{i})$. In a machine learning context, an action is also referred to as a target or label. Standard supervised learning methods can then be applied to $\mathcal{D}$ to train a model that predicts expert-like actions from observed states.
Imitation learning offers a data-driven approach for solving complex SDPs with combinatorial action spaces. A major challenge is the acquisition of high-quality expert actions, as optimal solutions are often computationally expensive. Beyond this, IL must scale to high-dimensional decision spaces and address covariate shift, where the learned policy encounters states during deployment that differ from those seen during training, leading to poor generalisation. This section introduces two IL paradigms, reviews key challenges, and discusses the use of noisy or imperfect experts as alternatives to expensive optimal experts.
3.1 Learning Paradigms
We describe two learning paradigms within IL: behavioural cloning and inverse reinforcement learning, followed by practical challenges and common remedies.
3.1.1 Behavioural Cloning (BC)
Behavioural cloning maps states, observations, or contexts to actions using expert demonstrations. The objective is to learn parameters $\theta$ by solving
$\theta=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{(x,a^{*})\in\mathcal{D}}\big[\mathcal{L}(x,a^{*},\pi_{\theta})\big].$ (5)
Equation˜5 minimises the one-step deviation error along expert trajectories. Within CO, two prominent approaches are common.
  • Parameter Prediction Task (PPT) - An ML model predicts parameters that define a CO problem, which is then solved by an external optimiser. Regret (also referred to as task loss in [donti2017task]) is typically defined as the difference between the optimal objective value with the knowledge of $\xi$ and the objective value realised by the prescriptive decision. Given state $x$ and a predictive model $\varphi_{\theta}$, we define the regret loss following [sadana2025survey] as
    $\mathcal{L}_{R}(x,a^{*},\theta)=C(x,a,\xi)-C(x,a^{*},\xi),$ (6)
    where $a=g(\varphi_{\theta}(x))$, $C$ denotes the total cost, and $g$ represents the optimisation solver. The loss in Equation˜6 measures the suboptimality of the induced decision $a$ relative to the expert action $a^{*}$ under realisation $\xi$. Training typically requires embedding a CO solver within the learning loop, often through a differentiable or problem-specific surrogate. The first part of the right-hand side of (6) ensures that the computed gap always considers the difference between the best achievable $a$ under $\theta$ and the optimal decision $a^{*}$. This gap always remains positive unless $a$ matches $a^{*}$, where the gap is zero. Hence, optimising this loss allows us to adjust $\theta$, to ensure that the CO-layer solution converges to the optimal solution. PPT is also referred to as the regret minimisation task (RMT) [sadana2025survey].
  • Action Imitation Task (AIT) - In this setting, a model $\pi_{\theta}$ directly outputs an action $a$ that approximates the expert action $a^{*}$, thereby reducing dependence on an external solver at inference time. A common imitation loss is
    $\mathcal{L}_{I}(x,a^{*},\theta)=\ell(a,a^{*}),$ (7)
    where $a=\pi_{\theta}(x)$ and $\ell(\cdot,\cdot)$ may be mean squared error or binary cross-entropy. When the regret objective is used as $\ell$, the approach aligns closely with PPT. AIT is also referred to as learning to optimise (LtO) [kotary2024learning].
[…]
Figure 1: PPT or RMT : Loss = $C({x,a},\boldsymbol{\xi})-C({x,a^{*}},\boldsymbol{\xi})$
$\boldsymbol{x}$$\min_{a\in\mathcal{F}}C(x,a,\bar{\xi})$$\pi_{\theta}$$\boldsymbol{\hat{a}}$  $\boldsymbol{\hat{a}}$Refer to captionRefer to caption Optimization PhasePrediction Phase
Figure 2: AIT or LtO: Loss = $\ell({a},{a^{*}})$
Comparing PPT and AIT
Under the PPT paradigm (Figure˜1), the ML model predicts parameters that define a combinatorial optimisation problem, which is then solved by an external optimiser. In contrast, the AIT framework (Figure˜2) eliminates this two-stage inference process by training the ML model to directly output a decision or action, thereby avoiding the need for a downstream optimisation solver at inference time. It is important to note, however, that an optimisation solver remains necessary during data generation to compute the expert actions $a^{*}$.
Imitation learning with AIT models in CO settings presents several challenges, primarily due to large and complex decision spaces and the presence of ambiguous demonstrations. Without explicitly structured output spaces, standard models may generate infeasible or inconsistent actions. Feasibility can be restored through post-processing techniques such as projection or heuristic correction, although this may result in a modest increase in regret [kotary2024learning]. The PPT approach avoids infeasibility by construction, as feasibility is enforced by the optimisation solver following parameter prediction [greif2024combinatorial, baty2024combinatorial]. This advantage comes at the cost of tighter coupling between the learning model and the solver, since the loss depends on the optimisation output and often requires differentiation through the solver. Such differentiation is typically complex and problem-specific, although gradient-based approaches remain popular due to their compatibility with neural network architectures [mandi2024decision, sadana2025survey]. Both PPT and AIT offer promising avenues for decision making in CO problems under uncertainty, yet practical understanding of their relative strengths remains limited.
Distribution Shift
Given an offline training dataset $\mathcal{D}=\{(x_{i},a_{i}^{*})\}_{i=1}^{N}$ generated by an expert policy, that is $(x_{i},a_{i}^{*})\sim\pi^{*}$, behavioural cloning seeks to solve Equation˜5. This formulation gives rise to a distribution shift, or distribution mismatch, a phenomenon well documented in the supervised learning literature for sequential decision-making problems. During deployment, deviations from expert behaviour may lead the learned policy to encounter states that were absent from the training data. Such deviations can compound over time, causing the policy to drift into unfamiliar regions of the state space where reliable recovery is difficult.
To achieve stronger generalisation, one would instead aim to solve
$\theta=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{x\sim\pi_{\theta}}\big[\mathcal{L}(x,a^{*},\pi_{\theta})\big],$ (8)
thereby training the model under the state distribution induced by its own actions. This formulation exposes the learner to states arising from suboptimal decisions and enables exploration of a broader portion of the transition dynamics, improving robustness at deployment.
A standard methodology for addressing this mismatch is dataset aggregation (DAgger), introduced by [ross2011reduction]. DAgger iteratively generates trajectories using the current learned policy and relabels the visited states with actions provided by the expert. This process aligns the training distribution with the deployment distribution. However, in multi-stage CO settings, DAgger can be computationally expensive, as expert labelling requires repeatedly solving optimisation problems at many decision epochs. As a result, DAgger involves intermittent but costly expert feedback throughout the training process.
3.1.2 Inverse Reinforcement Learning (IRL)
Inverse reinforcement learning (IRL), first proposed in [russell1998learning], seeks to recover the underlying reward function from expert demonstrations rather than learning a direct mapping from states to actions. Consistent with the IRL literature, we use reward-based notation, in contrast to the cost-based formulation adopted for behavioural cloning. Once the reward function is inferred, a policy can be learned by solving the resulting reinforcement learning problem.
The objective is to recover a reward function parameterised by $\theta$ such that the induced optimal policy matches the expert policy $\pi^{*}$, that is,
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}\Bigg[\sum_{t=1}^{T}r_{\theta}(x_{t},\pi(x_{t}))\Bigg].$ (9)
Given an expert dataset $\mathcal{D}$, the reward parameters are obtained by solving
$\theta=\operatorname*{arg\,max}_{\theta}\;\mathcal{J}(\mathcal{D},r_{\theta}).$ (10)
Reinforcement learning agents improve their behaviour through repeated interaction with the environment, allowing them to account for the long-term consequences of actions and recover more effectively from mistakes. As a result, IRL methods, which rely on RL for policy optimisation, are generally less sensitive to covariate shift than behavioural cloning. However, IRL is computationally demanding, as it requires repeated interaction with the environment and alternates between reward estimation and policy learning. IRL is also sample inefficient and inherently ill posed, since multiple reward functions can explain the same expert behaviour, making unique recovery of the true reward difficult.
Several classes of IRL methods have been proposed, differing mainly in how the objective function $\mathcal{J}$ is defined and optimised. Prominent approaches include maximum margin methods, which maximise the separation between expert and non-expert trajectories; maximum entropy IRL, which matches expert behaviour while encouraging stochasticity; Bayesian IRL, which infers a distribution over reward functions; and generative adversarial imitation learning (GAIL), which frames learning as a game between a policy and a discriminator. Comprehensive surveys of these methods are provided in [zare2024survey] and [zheng2022imitation].
IRL is particularly well-suited to settings where expert demonstrations are limited and accurate modelling of long-term behaviour is essential. A key advantage is that once the reward function is learned, policies can be re-optimised if system dynamics change, without requiring additional demonstrations. By contrast, behavioural cloning is more appropriate when expert data are abundant and cover the relevant state–action space, and when small deviations from expert behaviour do not lead to severe performance degradation.
3.2 Learning from Noisy or Suboptimal Experts
Many IL tasks rely on expert demonstrations, yet high-quality and fully optimal data are often difficult to obtain in practice. Human experts may make mistakes, and computing optimal solutions can be prohibitively expensive. Consequently, most real-world IL applications rely on imperfect but readily available demonstrations. In combinatorial optimisation (CO), human experts are typically unsuitable due to problem complexity, and optimisation solvers are used instead. However, these solvers are computationally demanding, particularly for sequential decision problems under uncertainty [powell2022reinforcement]. To mitigate this cost, demonstrations are often generated using static approximations or simplified scenarios. For example, [pham2023prediction, baty2024combinatorial] solve static CO problems offline to produce training targets. Although these solutions are not fully optimal, they are informative and substantially more tractable.
In such settings, suboptimal demonstrations are frequently treated as if they were optimal, and classical IL methods are applied directly [pham2023prediction, baty2024combinatorial]. However, given the $\mathcal{NP}$-hard nature of many CO problems, even static instances often require approximations. Solvers are commonly run with time limits, heuristics, or prescribed optimality gaps, all of which introduce additional noise into the demonstrations. Despite this prevalence, there is limited understanding of how such imperfections affect IL performance in CO settings, and systematic comparative studies remain scarce.
Several recent works explicitly consider learning under noisy supervision and adapt IL algorithms accordingly. [xu2022discriminator, kim2021demodice] propose offline IL frameworks that leverage supplementary datasets generated by suboptimal policies to enhance learning. [sekhari2024selective] introduces an IL algorithm that achieves low regret while reducing expert queries, even in the presence of noisy demonstrations. [wang2023imitation] presents a two-stage approach that first purifies noisy demonstrations using a diffusion process before applying imitation learning. Additional studies on learning from imperfect demonstrations include [wu2019imitation, sasaki2020behavioral, wang2021learning], as surveyed in [zare2024survey].
Another challenge arises when multiple labels are available for the same state, often generated by different noisy experts. For instance, drivers may follow different policies under identical conditions, resulting in conflicting demonstrations. While such diversity can enrich the learning signal, it necessitates aggregation or filtering mechanisms to prevent degradation in policy performance. Early work by [raykar2009supervised, raykar2010learning] examines this issue in pre-labelled datasets. More recently, [sun2023mega] propose MEGA-DAgger, an interactive learning framework designed for multiple imperfect experts, which filters unsafe demonstrations and evaluates experts using scenario-specific criteria. [sekhari2024selective] further extend their approach to multi-expert settings in which each expert specialises in a distinct region of the state space.
3.3 Applications of IL in SDPs
Recent research on applying machine learning to SDPs in combinatorial optimisation largely follows two complementary directions. The first direction focuses on ML-augmented optimisation, where learning methods are used to enhance the efficiency of classical optimisation algorithms such as branch-and-bound or column generation. The second direction, referred to as end-to-end learning (E2EL), integrates combinatorial solvers or optimisation layers directly within learning architectures. This classification follows the taxonomy proposed by [kotary2021end]. While both IL and RL are employed across these approaches, the discussion here emphasises recent work that applies IL to decision making within these two categories.
3.3.1 ML-Augmented Optimisation
ML-augmented optimisation incorporates machine learning models into the decision-making processes of optimisation algorithms. In this paradigm, ML may act as a preprocessing tool to configure or initialise an algorithm, or as a component invoked repeatedly during execution to guide key decisions [bengio2021machine]. This approach has been explored across a wide range of problem domains. Some methods focus on predicting discrete actions, such as branching or node selection decisions, while others predict scores or surrogate quantities associated with these actions.
  1. 1.
    Branch-and-Bound - We refer the reader to [bengio2021machine] and [scavuzzo2024machine] for a basic introduction to Branch-and-Bound (B&B). Branch-and-Bound (B&B) relies on two core decisions: selecting a branching variable and choosing a node to expand. Several studies use IL to guide these decisions. Works such as [alvarez2017machine, alvarez2016online, balcan2018learning, ding2020accelerating, gasse2019exact, gupta2022lookback, khalil2016learning, lin2022learning, lu-2020, Nair2020SolvingMI, zarpellon2021parameterizing] apply IL to learn branching policies. In parallel, [he2014learning, khalil2022mip, labassi2022learning, yilmaz2021study] focus on node selection strategies. The DAgger algorithm is employed in [he2014learning] and [Nair2020SolvingMI] to incorporate expert interaction. Additional contributions include learning when to invoke primal heuristics [khalil2017learning] and predicting backdoors to accelerate mixed-integer programming [cai2024multi].
  2. 2.
    Cutting Planes. Cut generation is a central technique in mixed-integer programming, aimed at tightening relaxations without excluding feasible integer solutions. Several studies use supervised learning to decide which cuts to add, including [deza2023machine, ding2020accelerating, huang2022learning, paulus2022learning].
  3. 3.
    Column Generation and Branch-and-Price. Column generation solves large-scale linear programmes by generating variables on demand. When combined with Branch-and-Bound, this yields the Branch-and-Price framework [barnhart1998branch]. Learning-based methods in this setting focus on selecting promising columns or designing effective branching strategies to improve speed. Representative works include [furian2021machine, morabit2021machine, pereira2022learning, sun2022learning, shen2022enhancing].
  4. 4.
    Large Neighbourhood Search. Large Neighbourhood Search (LNS) improves solutions by iteratively destroying and repairing partial solutions. Machine learning has been used to generate high-quality initial solutions and to adaptively select neighbourhoods [hizir2025large, huang2023searching, liu2022learning, mossina2019multi, Nair2020SolvingMI, song2020general, sonnerat2021learning].
  5. 5.
    Search Space Reduction. Another line of work aims to directly reduce the search space by eliminating decisions unlikely to appear in an optimal solution. For example, in the travelling salesman problem, ML models can predict irrelevant arcs for pruning. Studies following this approach include [fischetti2019machine, fitzpatrick2021learning, grassia2019learning, lauri2019fine, larocca2024combining, morabit2022machine, spieckermann2025reduce, sun2021generalization, tahir2021improved, tayebi2022learning, vaclavik2018accelerating].
  6. 6.
    Constraint Programming. Constraint Programming (CP) solves combinatorial problems through systematic constraint propagation and search. Several works leverage supervised learning to guide CP solvers by predicting constraint satisfaction or search decisions, including [selsam2018learning, sun2024enhancing, xu2018towards].
  7. 7.
    Algorithm Configuration. Machine learning is also used to decide whether specific algorithmic components should be applied to a given instance. Examples include predicting when to apply Dantzig–Wolfe decomposition [kruber2017learning], when to use scaling techniques [berthold2021learning], whether to apply cutting planes beyond the root node [berthold2206learning], and whether to linearise the objective in mixed-integer quadratic programmes [bonami2018learning]. In all cases, algorithm runs are used to generate labelled training data.
Problems in this category may be either single-stage or sequential. Single-stage settings include tasks such as predicting high-quality initial solutions for warm-starting. In contrast, sequential settings require a series of interdependent decisions, such as branching choices along a B&B tree. [Nair2020SolvingMI] addresses both perspectives. In addition, some approaches rely on a single prediction to inform multiple downstream decisions. For example, [khalil2022mip] predicts likely variable assignments at the root node of a binary linear program and reuses this information for subsequent node selection.
3.3.2 End-to-End Learning
End-to-end learning refers to a modelling paradigm that tightly integrates machine learning and optimisation within a single framework. In this setting, the learning model either directly predicts solutions to optimisation problems, corresponding to the AIT, or produces partially specified optimisation models whose parameters are inferred from raw or structured input data, corresponding to the PPT.
Parameter Prediction Task
Early work in this direction includes [donti2017task] and [elmachtoub2022smart]. In particular, [elmachtoub2022smart] introduce the Smart “Predict, then Optimise” (SPO) framework, which trains ML models to predict uncertain coefficients in linear objective functions. As the original SPO loss is non-convex and difficult to optimise, the authors propose a convex surrogate, SPO+, which enables efficient training. Subsequent studies apply the SPO and SPO+ losses across a range of applications, including last-mile delivery, energy systems, and ship management [chu2023data, demirovic2020dynamic, sang2022electricity, tian2023smart].
[donti2017task] extend regret minimisation to stochastic convex optimisation problems and rely on implicit differentiation of the regret objective. More recently, [baty2024combinatorial, greif2024combinatorial, jungel2025learning, rautenstrauss2025optimization] propose learning pipelines enriched with optimisation layers and apply them to dynamic vehicle routing problems. These works typically rely on full-information offline solutions as supervision and aim to minimise regret with respect to these benchmarks. Notably, [greif2024combinatorial] and [rautenstrauss2025optimization] employ DAgger-based data collection strategies to improve robustness.
Related ideas appear in [parmentier2021learning], which studies a single-machine scheduling problem using structured prediction [nowozin2011structured] to approximate hard instances via easier surrogate problems whose parameters are learned. [sinclair2023hindsight] leverage hindsight or full-information solutions to estimate expert values and propose a hindsight learning algorithm, demonstrating its effectiveness across multiple problem domains.
Action Imitation
An early attempt to directly predict solutions to difficult combinatorial problems, such as the Travelling Salesman Problem (TSP) is due to [vinyals2015pointer], who propose pointer networks trained in a supervised manner on solved instances. Their approach demonstrates limited generalisation to instances of varying sizes. Building on this idea, [li2018combinatorial] and [joshi2019efficient] develop supervised models for static combinatorial optimisation problems.
[kotary2024learning] propose a supervised learning framework for predicting optimal solutions based on observable features correlated with unknown parameters of a parametric optimisation problem. [chen2023two] address the Flexible Job Shop Scheduling Problem using a two-stage supervised framework that imitates a constraint programming solver, first predicting machine assignments and then job start times, followed by post-hoc feasibility restoration.
Several works extend action imitation to dynamic or stochastic settings. [larsen2018predicting] study the Load Planning Problem by imitating second-stage decisions of a two-stage optimisation model using full-information supervision. [ozarik2024machine] learn driver routing behaviour in last-mile delivery from historical data. [qi2023practical] consider dynamic inventory management under demand uncertainty, training supervised models on offline full-information solutions. [pham2023prediction] apply regression-based learning to dynamic radiotherapy appointment scheduling. More recently, [gawas2024ildagger] study the Employee Notification Timing Problem and use DAgger to train policies that imitate a range of offline expert models.
ML-augmented optimisation and end-to-end learning thus represent two dominant paradigms for combining machine learning with optimisation. While ML-augmented approaches embed learning within specific algorithmic components such as branching, cutting, or search space reduction, end-to-end methods aim to predict decisions or optimisation parameters directly through unified learning pipelines. Across both paradigms, supervision plays a central role, yet the nature of the expert varies substantially, ranging from full-information solutions and heuristics to intermediate solver states. These choices have significant implications for learning performance, generalisability, and computational cost. Understanding how experts are defined and utilised is therefore crucial for algorithm design. We next introduce a taxonomy of expert types to clarify these distinctions and to guide the development of imitation learning methods for sequential decision problems.
4 Expert Taxonomy
An expert in imitation learning is a policy, or more generally a decision-making function, that specifies the desired or target behaviour in a sequential decision problem. The expert guides the learning agent by demonstrating actions in given states, typically under the assumption of access to an optimal or near-optimal strategy with respect to a known or implicit objective. Alternative terms used in the literature include oracle, demonstrator, optimal policy, ground truth, reference policy, and target. Broadly, we distinguish between two categories of experts.
  • Human Expert. Human experts rely on practical experience and contextual understanding to make decisions under uncertainty. In dynamic environments such as vehicle dispatching or healthcare scheduling, they implicitly account for real-time information, operational constraints, and institutional knowledge. For example, [ozarik2024machine] study last-mile delivery and imitate driver decisions that capture practical insights not explicitly modelled in traditional optimisation frameworks. While such decisions often reflect nuanced judgement, they may also be inconsistent or influenced by cognitive biases. Moreover, collecting sufficiently large and high-quality datasets from human experts can be challenging, particularly in domains where expert time is scarce or costly.
  • Algorithmic Expert. Algorithmic experts generate decisions using mathematical models, optimisation solvers, or rule-based and heuristic procedures. These experts can typically produce large volumes of consistent and high-quality demonstrations. Their structure depends on how the decision objective is defined and how uncertainty, denoted by $\boldsymbol{\xi}$, is treated. The vector $\boldsymbol{\xi}$ represents exogenous factors such as demand, travel times, or arrival processes, which are not fully known at the time decisions are made.
Given the complexity of real-world SDPs and the need for timely decisions, model-based algorithmic experts are often more suitable for generating demonstrations in practice.
Ideally, for a sequential decision problem, the optimal action at decision epoch $k$ is obtained by solving
$a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\mathbb{E}\Big\{C_{k}(x_{k},a_{k})+V_{k+1}(x_{k+1})\,\big|\,x_{k}\Big\},$ (11)
where the expectation is taken with respect to future uncertainty. For many real-world business applications, computing such optimal actions, even offline, is often impractical due to problem complexity. Consequently, it is common in operations research to design algorithms that produce near-optimal solutions—decisions that are not globally optimal but are sufficiently close to be useful in practice. When such solutions can be computed within reasonable time limits, they can serve as expert demonstrations for IL.
Despite the widespread use of approximate or surrogate experts, the literature lacks a clear and systematic taxonomy for defining experts and their roles within IL. To address this gap, we propose a structured taxonomy of expert types to guide the design and application of IL methods in sequential decision problems.
4.1 Expert Decision Models under Uncertainty
We further classify algorithmic experts according to how they handle uncertainty $\boldsymbol{\xi}$, which captures factors such as future demand, travel times, or patient arrivals that are not fully observed at the time of decision making. The treatment of $\boldsymbol{\xi}$ determines whether an expert behaves myopically or anticipates future outcomes, and whether it is relatively simple or highly adaptive. We outline five classes of algorithmic experts, distinguished by how uncertainty is incorporated into their decision models.
  1. 1.
    Myopic Expert - A myopic expert selects actions based solely on immediate costs, effectively ignoring future uncertainty. It treats $\boldsymbol{\xi}$ as unavailable or irrelevant and bases decisions only on the current state. This leads to simple models and fast solution times, but often results in poor long-term performance due to the lack of anticipation.
  2. 2.
    Deterministic Expert - The deterministic expert replaces uncertain parameters $\boldsymbol{\xi}$ with fixed point estimates $\hat{\boldsymbol{\xi}}$, such as mean or nominal values, and solves the resulting deterministic optimisation problem. While straightforward to implement, this approach may perform poorly when actual realisations deviate substantially from the assumed estimates.
  3. 3.
    Full-Information (Hindsight) Expert - The full-information expert assumes perfect foresight and solves the optimisation problem using the realised values $\tilde{\boldsymbol{\xi}}$ of all uncertain parameters. Although infeasible in real-world operations, this expert serves as a useful benchmark, as it yields the best possible decision in hindsight for each instance.
  4. 4.
    Two-Stage Stochastic Expert - In the two-stage stochastic framework, uncertainty $\boldsymbol{\xi}$ is not observed at the initial decision stage. The expert first selects a here-and-now action and subsequently adapts through recourse decisions once uncertainty is revealed. This structure captures limited anticipation while retaining tractability and is widely used in stochastic programming.
  5. 5.
    Multi-Stage Stochastic Expert - Multi-stage stochastic experts operate in fully dynamic settings where uncertainty unfolds gradually over time. Decisions are updated at each stage based on newly revealed information, allowing policies to adapt as $\boldsymbol{\xi}$ evolves. While this formulation closely reflects real-world decision making under uncertainty, it is computationally expensive to solve exactly, particularly when uncertainty is high-dimensional or difficult to model.
Table 1 summarises the objective structure and treatment of uncertainty for each expert type. Here, $C$ denotes the total cost function and $\mathcal{A}$ represents the set of all feasible actions.
Policy/Action Model Type Objective Uncertainty
$\pi^{*m},a^{*m}$ Myopic $\operatorname*{arg\,min}\limits_{a\in\mathcal{A}}\{C(a)\}$ No uncertainty
[…]
$\pi^{*t},a^{*t}$ Two-Stage Stochastic $\operatorname*{arg\,min}\limits_{a\in\mathcal{A}^{1}}\{C(a,\xi)=Q_{1}(a)+\mathbb{E}_{\xi}[Q_{2}(a,\xi)]\}$ Uncertainty Set
$\pi^{*},a^{*}$ Multi-Stage Stochastic $\operatorname*{arg\,min}\limits_{a\in\mathcal{A}}\mathbb{E}_{\xi}\{C(a,\xi)\}$ Full Uncertainty
Table 1: Summary of Expert Models
Most of the literature in SDP relies on either myopic or full-information experts for imitation, depending on the problem structure and availability of future information. In settings where uncertainty is absent—such as static or fully observable optimisation problems—a deterministic expert is typically employed. Since all relevant parameters are known with certainty, the expert can compute a single optimal solution using standard optimisation techniques.
4.2 Expert Optimality
Even when employing the expert types defined above, the underlying optimisation problem may remain difficult to solve. Many deterministic optimisation problems are $\mathcal{NP}$-hard, rendering exact solution approaches computationally impractical in many real-world settings. In such cases, practitioners rely on heuristics or impose stopping criteria when using optimisation solvers. We therefore classify experts according to the optimality guarantees of the models they solve.
  • Task-Optimal Expert. A task-optimal expert solves its underlying optimisation model to optimality, assuming the problem is tractable and can be solved exactly within reasonable computational limits. Such experts are most applicable when the problem structure permits efficient solution methods. For example, a strong branching expert evaluates linear programming relaxations corresponding to all fractional variables and selects the branching variable that yields the greatest improvement in the objective [Nair2020SolvingMI, alvarez2016online]. Optimal solutions may be obtained using a variety of specialised solvers, including mixed-integer programming solvers [sun2022learning], SAT solvers [sun2024enhancing], and dedicated solvers such as Concorde for the travelling salesman problem [sun2021generalization]. We denote the policy of a task-optimal expert by $\pi^{*\cdot}$ and the corresponding optimal action by $a^{*\cdot}$ where $\cdot$ would denote the superscript of the respective experts from Table˜1.
  • Approximate Expert. An approximate expert produces actions that aim to closely approximate optimal or near-optimal decisions while trading off solution quality for computational efficiency. Several commonly used approximation strategies include:
    • Time limits - The solver is terminated after a fixed time budget, returning the best solution found so far [li2025learning, baty2024combinatorial, pham2023prediction, sun2021generalization].
    • Optimality gaps -The solver terminates once a predefined optimality gap is achieved, ensuring a bounded deviation from optimality [larsen2018predicting].
    • Objective-based refinement - Alternative formulations or local improvements are used to incrementally improve a solution relative to the objective function [ding2020accelerating].
    • Heuristics - High-quality heuristic methods may be used directly to generate solutions. For example, [baty2024combinatorial] employ hybrid genetic search for the prize-collecting TSP. Alternatively, a finite set of candidate actions may be evaluated, and the best-performing action selected [song2020general].
    We denote the policy of an approximate expert by $\pi^{\cdot}$ and the corresponding action by $a^{\cdot}$.
4.3 Expert Interaction
Depending on the cost, availability, and reliability of the expert, different strategies can be used to incorporate expert supervision into IL.
  • Learning without interaction - In this non-iterative setting, the expert is queried once to label a batch of observed states, after which the learning model is trained offline. This approach is appropriate when expert queries are expensive or time-consuming, as it minimises interaction. Most studies in the literature adopt this strategy by generating a fixed dataset of expert-labelled states. However, as discussed in Section˜3, one-shot learning is vulnerable to poor generalisation when the learned policy encounters states that were not present in the expert dataset. We also refer to this setting as Isolated learning.
  • Learning with interaction - Interactive expert querying is commonly used to enrich training data and mitigate distribution shift by exposing the learner to a broader range of states. [he2014learning] employ the DAgger algorithm to iteratively query an expert when learning node selection policies for branch-and-bound. Similarly, [song2020general] propose a DAgger-based approach to learn search policies for both $A^{*}$ and branch-and-bound algorithms in integer programming. Several recent works adopt iterative expert interaction in combinatorial optimisation settings, including [gawas2024ildagger, greif2024combinatorial, Nair2020SolvingMI, rautenstrauss2025optimization].
    While effective, the standard DAgger algorithm can be computationally expensive due to frequent expert queries. To reduce this cost, alternative interaction schemes have been proposed [labassi2022learning, rautenstrauss2025optimization]. Related work also considers online learning settings, where models are updated incrementally as new data are collected, which likewise requires expert interaction. Examples include [alvarez2016online, khalil2016learning, gasse2019exact, lu-2020].
Figure˜3 summarises the literature on expert models, highlighting distinctions between task-optimal and approximate experts, as well as between one-shot and interactive interaction strategies. The figure indicates that stochastic expert models and interactive learning approaches remain relatively uncommon, likely due to the additional methodological and computational complexity they introduce.
LiteratureMyopicDeterministicHindsight2-Stage Stochastic[alvarez2017machine], [furian2021machine], [gupta2022lookback],[khalil2017learning], [kraul2023machine], [morabit2021machine],[Nair2020SolvingMI], [ojha2023optimization], [paulus2022learning],[shen2022enhancing], [sonnerat2021learning], [sun2022learning][gasse2019exact], [khalil2016learning], [vaclavik2018accelerating][guaje2024machine], [huang2023searching], [lin2022learning],[paul2025data], [pereira2022learning], [zarpellon2021parameterizing][lu-2020][demirovic2020dynamic], [elmachtoub2022smart], [julien2024machine],[jungel2025learning], [kotary2024learning], [niroumandrad2024learning],[qi2023practical], [sinclair2023hindsight], [wang2023imitation][he2014learning], [gawas2024ildagger], [labassi2022learning],[rautenstrauss2025optimization][baty2024combinatorial], [cai2024multi], [chu2023data],[donti2017task], [hottung2020deep], [larsen2018predicting],[li2025learning], [pham2023prediction], [sang2022electricity],[tian2023smart], [yilmaz2021study][song2020general][detassis2021teaching], [fitzpatrick2021learning], [joshi2019efficient],[kaempfer2018learning], [lauri2019fine], [li2018combinatorial],[liu2022learning], [lodi2020learning], [morabit2022machine],[parmentier2021learning], [selsam2018learning], [spieckermann2025reduce],[sun2021generalization][gawas2024ildagger] [chen2023two], [ding2020accelerating], [gerbaux2025machine],[grassia2019learning], [han2023gnn], [hizir2025large],[khalil2022mip], [larocca2024combining], [matsuoka2019machine],[mossina2019multi], [tayebi2022learning], [vinyals2015pointer],[xavier2021learning], [xu2018towards], [yaakoubi2020flight],[yang2023learning][greif2024combinatorial],[Nair2020SolvingMI][abbasi2020predicting][gawas2024ildagger]OptimalApproximateOne-ShotInteractive
Figure 3: Classification of literature.
4.4 Labelling strategy
Most existing studies rely on an expert to provide a single label for each state. However, for certain types of experts, it is possible to obtain multiple labels for the same state within one or more interactions.
  • Single label per state - This is the standard approach, in which the expert is queried once to obtain a single target action for a given state.
  • Multiple labels per state - Expert actions may vary due to approximation methods or inherent stochasticity in the underlying models. To improve robustness, multiple labels can be collected for the same state and combined using aggregation techniques such as majority voting or averaging. This process reduces variability and mitigates the impact of expert error. Multiple labels can be obtained using two mechanisms.
    1. 1.
      Single call - A single interaction with the expert yields multiple candidate actions for a given state. This occurs, for example, when solving an integer programme using a mathematical programming solver that returns multiple feasible solutions. In [khalil2022mip], CPLEX is allowed to run for 60 minutes to generate a solution pool for each instance, terminating after identifying 1,000 solutions with objective values within 10% of the best solution found.
    2. 2.
      Multiple calls - Alternatively, repeated queries to the expert may be used to collect several actions for the same state. This approach is common when using deterministic experts in dynamic stochastic programming, where the optimal action depends on sampled future scenarios [gawas2024ildagger, greif2024combinatorial].
4.5 Number of Experts Used
Algorithms can also be categorised based on the number of experts involved and how their expertise is integrated during learning. We identify two primary categories:
  • Single Expert- The conventional setting in imitation learning, where a single, typically optimal expert provides guidance. The learner assumes the expert is both reliable and inexpensive to query, consistently supplying accurate actions for all states. This simplifies learning but may limit adaptability in complex or diverse environments. Most literature uses a single expert to build the dataset.
  • Multi-Expert- In some cases, multiple experts with complementary specialisations are available. Their actions can be aggregated into a composite "meta-expert", combining diverse expertise to enhance robustness and flexibility across different aspects of the task. [song2020general] uses two different experts to get demonstrations. First, they use an expensive expert to gather initial expert demonstrations, which are used to train initial imitation learning policies. Next, a simpler retrospective expert is used to augment the dataset. [rautenstrauss2025optimization] uses a set of non-optimal policies to label states for which the true hindsight expert is expensive. These states generally represent initial decision epochs in the horizon. [sun2023mega] introduces MEGA-DAgger, a variant of DAgger designed for interactive learning with multiple imperfect experts.
4.6 Trajectory Control
When experts are employed within an iterative learning framework, they can also be used to control trajectories and correct errors made by a previously learned model. This approach ensures that diverse states are seen that are likely to be encountered during deployment. The following control strategies are commonly used:
  • Total Expert Control- The expert governs all states visited within a learning trajectory. A typical example is pure supervised learning without iteration, where the expert labels all states once without further interaction.
  • Diminishing Expert Control- The learner initially relies heavily on the expert, but gradually reduces this dependence as its performance improves. This approach is exemplified by the basic DAgger algorithm, also referred to as Vanilla DAgger [ross2011reduction]. This type of control is most used in interactive experts in [gawas2024ildagger], [greif2024combinatorial], [song2020general].
  • Conditional Expert Control- The expert is invoked selectively based on performance criteria, such as the confidence of the learned model action falling below a threshold, as in SAFE DAgger [zhang2016query]. Alternatively, an external policy may guide trajectories when the true expert is expensive or imperfect.
Moreover, to reduce reliance on costly expert calls, the expert is queried only for critical cases, such as novel or poorly handled states. This selective querying improves sample efficiency and focuses learning on high-impact scenarios. [menda2019ensembledagger], [hoque2021lazydagger] describe the Ensemble DAgger and Lazy DAgger algorithms, where the expert is selectively queried based on the discrepancy measure between the expert and learned policy actions. However, such variants of the DAgger algorithms have not been experimented with for SDP in operations research.
Using the expert taxonomy, we now propose and test a general DAgger-based algorithm that links each dimension of the taxonomy to a corresponding design choice in the learning procedure, enabling a structured comparison of expert configurations.
5 Problem Description
The classical Generalised Assignment Problem (GAP) has been extensively studied since the seminal work of [balachandran1976integer]. The problem consists of assigning a set of items (or jobs) to a set of agents (or machines) such that each job is assigned to exactly one agent, agent capacity constraints are respected, and the total assignment cost or profit is optimised. GAP has found applications in a wide range of domains, including job scheduling [balachandran1976integer], routing [fischetti2019machine], and patient admission scheduling [ceschia2011local]. A standard integer programming formulation of the GAP is given by
$\max\quad$ $\sum_{i=1}^{m}\sum_{j=1}^{n}c_{ij}x_{ij}$ (12)
s.t. $\sum_{i=1}^{m}x_{ij}=1,\quad$ $\forall j=1,\dots,n,$ (13)
$\sum_{j=1}^{n}d_{ij}x_{ij}\leq L_{i},\quad$ $\forall i=1,\dots,m,$ (14)
$x_{ij}\in\{0,1\},\quad$ $\forall i=1,\dots,m,\ \forall j=1,\dots,n.$ (15)
Here, $c_{ij}\in\mathbb{R}_{+}$ denotes the profit obtained by assigning job $j$ to agent $i$, $d_{ij}\in\mathbb{R}_{+}$ is the resource consumption of agent $i$ when processing job $j$, $L_{i}\in\mathbb{R}_{+}$ is the capacity of agent $i$, and $x_{ij}$ is a binary decision variable equal to $1$ if job $j$ is assigned to agent $i$ and 0 otherwise.
5.1 Physician-to-Patient Assignment
We consider a dynamic physician-to-patient assignment (PPA) problem as a specific application of the GAP. In a clinical setting, appointment requests arrive sequentially over a booking horizon until a fixed cut-off time. The clinic is staffed by a set of physicians providing services during a given session. Let $\mathcal{P}=\{1,\dots,P\}$ denote the set of physicians. Each physician $p\in\mathcal{P}$ has a regular session duration $T$ (in minutes) and a maximum appointment capacity $L_{p}$.
Appointment requests arrive independently over time and are handled individually by a receptionist. At each arrival, the receptionist must decide whether to accept or reject the request based on the current system state. If the request is accepted, the receptionist must assign the patient to a physician. Each patient $k$ has a preferred physician $p_{k}$ among a set of eligible physicians $\mathcal{P}_{k}\subseteq\mathcal{P}$. Each patient is characterised by a service duration $d_{k}\in\mathbb{R}_{+}$, which is estimated at the time of the call, and a priority class $r_{k}\in\{1,2\}$. The arrival process is such that call from priority 1 patients arrive near the end of the horizon.
Rejecting a patient of priority $r_{k}$ incurs a penalty $c^{\text{rej}}_{r_{k}}$, with $c^{\text{rej}}_{1}>c^{\text{rej}}_{2}$. In addition, assigning a patient to an undesired but eligible physician incurs a preference penalty $c^{\text{pref}}_{r_{k}}$. The primary source of uncertainty in the system arises from the random number and sequence of patient arrivals within a session. This creates a trade-off between accepting lower-priority patients early and preserving capacity for potential high-priority patients who may arrive later. Furthermore, assigning patients to undesired physicians may reduce future capacity, potentially forcing rejections later in the booking horizon.
The resulting problem is a sequential decision problem in which each decision epoch corresponds to a patient arrival, and the action is to assign the patient to an eligible physician or to reject the request. The offline mixed-integer programming formulation, denoted by $(\textbf{{MIP}}_{\textbf{PPA}})$, is provided in the appendix.
5.2 MDP Formulation
We now formulate the PPA problem as a Markov decision process (MDP).
Decision Epochs – Let $\mathcal{K}=\{0,1,\dots,K\}$ denote the set of decision epochs, where each epoch corresponds to the arrival of a patient call-in. Here, $K$ denotes the total number of requests within a session.
State Variable – At decision epoch $k$, the system state is given by
$x_{k}=\big(x^{k}_{\text{patient}},x^{k}_{\text{physician}},x^{k}_{\text{eligibility}}\big).$
The component $x^{k}_{\text{patient}}=[d_{k},r_{k}]$ represents the characteristics of the current patient. The component
$x^{k}_{\text{physician}}=[x^{k}_{p}]_{p\in\mathcal{P}}$
captures the status of all physicians, where
$x^{k}_{p}=\big(\mathbbm{1}_{p=p_{k}},x^{k}_{l},x^{k}_{p1},x^{k}_{pw}\big).$
Here, $\mathbbm{1}_{p=p_{k}}$ indicates whether physician $p$ is the patient preferred physician, $x^{k}_{l}$ denotes the remaining assignment capacity, $x^{k}_{p1}$ is the number of priority-1 patients already assigned, and $x^{k}_{pw}$ is the current workload. The eligibility vector is given by a list of indicator functions
$x^{k}_{\text{eligibility}}=[\mathbbm{1}_{p\in\mathcal{P}_{k}}]_{p\in\mathcal{P}}.$
Decision Variable – At each decision epoch $k$, the receptionist selects an action
$a_{k}\in\mathcal{A}_{k}=\{0\}\cup\mathcal{P}_{k},$
where $a_{k}=0$ corresponds to rejecting the request and $a_{k}=p$ corresponds to assigning the patient to physician $p$. A physician $p$ is eligible only if $d_{k}+x^{k}_{pw}\leq L_{p}$, ensuring feasibility of all actions in $\mathcal{A}_{k}$.
Exogenous Information – The exogenous information at epoch $k$ corresponds to the characteristics of the next arriving patient and is denoted by
$\xi_{k}=[d_{k+1},r_{k+1},p_{k+1},\mathcal{P}_{k+1}].$
Transition Function – The transition function $X^{M}(\cdot)$ maps the system from state $x_{k}$ to $x_{k+1}$ according to
$x_{k+1}=X^{M}(x_{k},a_{k},\xi_{k}).$
Specifically, for the selected physician $p$,
  • $x^{k+1}_{l}=x^{k}_{l}-\mathbbm{1}_{a_{k}=p}$,
  • $x^{k+1}_{p1}=x^{k}_{p1}+\mathbbm{1}_{a_{k}=p}\mathbbm{1}_{r_{k}=1}$,
  • $x^{k+1}_{pw}=x^{k}_{pw}+\mathbbm{1}_{a_{k}=p}d_{k}$.
Cost Function – At each decision epoch, a cost $C_{k}(x_{k},a_{k})$ is incurred, defined as
$C_{k}(x_{k},a_{k})=\mathbbm{1}_{a_{k}=0}\,c^{\text{rej}}_{r_{k}}+\mathbbm{1}_{a_{k}\neq 0}\,\mathbbm{1}_{a_{k}\neq p_{k}}\,c^{\text{pref}}_{r_{k}}.$
6 Algorithm
In this section, we present a general-purpose algorithm for learning from experts in sequential decision problems. The proposed algorithm, shown in Algorithm˜1, explicitly incorporates the expert taxonomy introduced in Section˜4. A summary of the algorithm parameters is provided in Table˜2.
The algorithm is based on the Dataset Aggregation (DAgger) framework, which follows an iterative policy learning paradigm inspired by online learning and assumes access to one or more experts. Let $\mathcal{E}$ denote the set of available experts. At each learning iteration, indexed by $i$, the learner retrains its policy using all states encountered so far, together with the corresponding expert-labelled actions. This iterative process allows the learner to correct earlier mistakes and progressively improve their performance.
The procedure starts from an initial policy $\pi_{0}$, which may be obtained from expert demonstrations or derived from a heuristic. This policy is then executed for $H$ simulated episodes using the decision rule, each consisting of $K$ decision epochs. At each epoch, the learner observes the current state and may query an expert up to $J$ times, as determined by the function $F_{j}$. Each query returns an expert action, potentially from different experts in $\mathcal{E}$, depending on the expert-selection function $F_{e}$.
We denote an expert action by $a^{\cdot}_{ijk}$, where $i$ indexes the learning iteration, $j$ the expert call within the epoch, and $k$ the decision epoch. For clarity, we omit the episode index in this notation. This may be task-optimal or approximate depending on the complexity of the underlying optimisation problem. The parameter $\kappa$ specifies the list of expert stopping criteria, thereby controlling the quality of the returned action. A decision rule (DR) then determines which action $a_{k}$ is to be executed in the environment, influencing the trajectory of the episode. The resulting state–action pairs are stored in the dataset before the transition to the next state occurs.
After completing $H$ episodes, the dataset is augmented with the newly collected expert-labelled samples, and a new policy $\pi_{i+1}$ is trained. This process is repeated until a stopping condition is satisfied, yielding progressively refined policies. The following paragraphs relate the components of the expert taxonomy to Algorithm˜1.
Parameter Description
$\mathcal{D}$ Dataset containing state–expert action pairs
$H$ Number of times to simulate current learned policy $\pi_{i}$
$\mathcal{E}=\{\pi^{*}_{1},\pi^{*}_{2},\dots\}$ Set of available experts
$K$ Episode trajectory length
$F_{j}$ Function that determines the number of expert calls at a given state
$J$ Number of expert calls in a given state
$F_{e}$ Function that selects which expert to query
$\kappa$ Stopping criterion for the expert
AA Aggregation function for combining multiple expert actions
DRDR Decision rule for selecting the executed action
STOP() Stopping criterion
Table 2: List of parameters for Algorithm˜1
1
2
3Function DAgger():
4   Initialize dataset $\mathcal{D}\leftarrow\emptyset$, initial policy $\pi_{0}$, and iteration counter $i\leftarrow 0$;
5
6 while STOP() do
    // Simulate $H$ episodes using current policy $\pi_{i}$
7
8    for $h\in\{1,\dots,H\}$ do
       // Sample a $K$-step trajectory
9
10       for $k\in\{1,\dots,K\}$ do
11
12           Observe state $x_{k}$;
13
14          $J\leftarrow F_{j}(x_{k})$;
          // Make $J$ expert calls
15          for $j\in\{1,\dots,J\}$ do
             // Select expert
16             $\pi^{*}\leftarrow F_{e}(x_{k},\mathcal{E},j)$;
17             ${a}^{\cdot}_{ijk}\leftarrow\pi^{*}(x_{k},\kappa)$;
18
19            end for
20
21          $\hat{a}_{ik}\leftarrow A\big(\{{a}^{\cdot}_{ijk}\}_{j=1}^{J}\big)$;
22
23           Append $(x_{k},\hat{a}_{ik})$ to dataset $\mathcal{D}$;
24
25          $a_{k}\leftarrow DR(x_{k},\hat{a}_{ik},\pi_{i})$;
26
27          $x_{k+1}\leftarrow X^{M}(x_{k},a_{k},\xi_{k})$;
28
29         end for
30
31      end for
32
33    $i\leftarrow i+1$;
34      Train policy $\pi_{i+1}$ on dataset $\mathcal{D}$;
35
36   end while
37
38 end
Algorithm 1 General DAgger Algorithm
1Function VanillaDAgger($x_{k},\hat{a}_{ik},\pi_{i}$):
2 $a_{k}\leftarrow\pi_{i}(x_{k})$;
3 $\beta_{i}\leftarrow\lambda^{i-1}\beta_{0}$;
4 $z\sim\text{Uniform}(0,1)$;
5
6 if $z\leq\beta_{i}$ then
7    return $\hat{a}_{ik}$;
8
9 else
10    return $a_{k}$;
11
12   end if
13
14 end
Algorithm 2 VANILLA DAgger Decision Rule
Expert Type.
The set of experts $\mathcal{E}$ may include different model types, such as Myopic, Deterministic, Full-Information, and Two-Stage Stochastic experts. Multi-stage stochastic experts are typically excluded due to their high computational cost.
Expert Optimality.
The optimality of expert actions depends on the complexity of the underlying SDP. As such problems often involve large combinatorial decision spaces, approximation techniques are commonly employed. For example, a Two-Stage Stochastic expert may be restricted to a limited number of scenarios to maintain tractability. The parameter $\kappa$ specifies the expert stopping criterion and may represent a time limit, a target optimality gap, or other termination conditions or even a combination of multiple criteria.
Expert Interaction.
Experts may be queried either once or iteratively. In a one-shot setting, expert interaction is limited to a single iteration ($I=1$), after which the policy is trained offline. In interactive settings, the stopping criterion allows the algorithm to perform $I$ iterations, resulting in a sequence of learned policies $\{\pi_{1},\dots,\pi_{I}\}$.
Labelling Strategy.
The function $F_{j}$ determines the number of expert calls $J$ made at a given state. It may also skip expert queries for selected states, for instance when querying is expensive or when certain states provide limited learning value. Selective querying strategies have been explored in the literature [zhang2016query]. If an expert returns multiple actions in a single call, then ${a}^{*}_{ijk}$ represents a vector of actions. The aggregation function $A$ combines these actions into a single target label.
Number of Experts.
When $|\mathcal{E}|>1$, the algorithm operates with multiple experts. The function $F_{e}$ determines which expert is queried for each of the $J$ calls at a given state.
Trajectory Control.
The decision rule (DR) specifies which action is executed at each decision epoch. Many ML-based approaches rely on full expert control, executing expert actions throughout the trajectory. A common alternative is the Vanilla DAgger rule shown in Algorithm˜2, which executes the expert action with probability $\beta_{i}$ and the learner’s action with probability $1-\beta_{i}$. The probability $\beta_{i}$ typically decreases over iterations as a function of $\lambda$ and $\beta_{0}$, allowing the learner to gradually transition from expert-guided behaviour to autonomous decision making.
7 Results
To demonstrate the practical effectiveness of Algorithm˜1, we evaluate its performance on the physician-to-patient assignment (PPA) problem. All expert optimisation models are solved using Gurobi. The full experimental codebase is publicly available at GitHub. All experiments are conducted on a machine equipped with an Intel Xeon Gold 6258R CPU running at 2.70 GHz. The remainder of this section is organised as follows. Section˜7.1 describes the experimental setup, including instance generation, evaluation metrics, baseline policies, and the computational environment. Section˜7.2 presents and discusses the experimental results in detail, while Section˜7.3 draws inferences from these results.
7.1 Experimental Setup
This subsection describes the instance generation procedure and the configuration of the learning models and baseline methods.
Data
We begin by describing the parameters used to generate instances of the PPA. The proposed approach is evaluated on synthetically generated instances constructed under controlled conditions. The total number of patient calls arriving at the clinic, denoted by $K$, is sampled from a normal distribution with mean $\mu=100$ and standard deviation $\sigma=8$. Each patient is assigned a priority class: class 1 (high priority) with probability $0.3$, and class 2 (regular priority) otherwise.
To capture the tendency of high-priority patients to call later in the day, we generate an arrival score for each patient from a beta distribution, $\mathrm{Beta}(a,b)$. Specifically, arrival scores are drawn with parameters $(a=3,b=1)$ for class 1 patients and $(a=1,b=1)$ for class 2 patients. Patients are then ordered according to these arrival scores, ensuring that high-priority patients are more likely to appear towards the end of the scheduling horizon.
Appointment durations are assumed to be known at the time of the call and are sampled from a lognormal distribution. For class 1 patients, durations are generated using parameters $\mu=3$ and $\sigma=0.8$, whereas for class 2 patients the parameters are $\mu=2.3$ and $\sigma=0.3$. We consider a setting with $P=4$ physicians. To construct each patient’s eligibility set, physician weights are first sampled from a Dirichlet distribution with parameter vector $[0.4,0.3,0.15,0.15]$. The top $k$ physicians with the largest weights are then selected, where $k$ is drawn uniformly from $\mathcal{U}\{1,\dots,P\}$. We describe this process in detail in the appendix.
Rejection penalties are set to $c^{\text{rej}}_{1}=200$ for priority 1 patients and $c^{\text{rej}}_{2}=50$ for priority 2 patients. Assigning a patient $i$ to a non-preferred but eligible physician incurs a penalty of $c^{\text{pref}}_{r_{i}}=0.1\times c^{\text{rej}}_{r_{i}}$.
ML Models
We construct a simulation model of the PPA using the data generation parameters described above. At each decision epoch, the learned policy is invoked to select an action. For each decision epoch, we generate a training example of the form $(x_{k},\hat{a}_{k})$, where $x_{k}$ denotes the concatenation of the state features and $\hat{a}_{k}$ is the target action provided by the expert at epoch $k$. The learning model maps the current state features to a decision that either assigns the patient to an eligible physician or rejects the request.
As the learning architecture, we employ a multilayer perceptron (MLP) consisting of an input layer, a single hidden layer with 64 neurons, and an output layer. Normalised state features are provided as input, and the MLP outputs a probability distribution over the available actions, namely assignment to each eligible physician or rejection. The final action is selected based on this predicted probability vector.
Baselines
We compare the learned policies against three baseline methods. The first is a greedy policy, denoted by $\pi_{\text{greedy}}$, which assigns an arriving patient to their preferred physician whenever possible. If the preferred physician is unavailable, the patient is randomly assigned to any eligible physician. This policy also serves as a myopic expert.
The second baseline, denoted by $\pi^{t}$, corresponds to a two-stage stochastic expert. At each decision epoch, this policy samples 30 future demand scenarios and invokes the expert solver to determine an action.
The final baseline is the exact optimal policy, denoted by $\pi^{*o}$, computed separately for each test instance. This policy represents an upper bound on achievable performance and serves as a benchmark for assessing the quality of learned and heuristic policies.
7.2 Performance Evaluation
We evaluate multiple configurations of expert models using Algorithm˜1 and compare their performance against the baseline policies. All methods are tested on a dataset of 1,000 instances. For each policy, we report the average objective value, the total number of rejected patients by priority class, and the total number of assignments to non-preferred physicians, aggregated across both priority classes. Each expert model is given a time limit of 30 seconds to compute an action at each decision epoch.
7.2.1 Baselines
Table˜3 summarises the performance of the baseline policies. The greedy policy $\pi_{\text{greedy}}$ performs poorly, primarily due to a high number of rejections among priority 1 patients. In contrast, the two-stage stochastic policy $\pi^{t}$ achieves substantially better performance by reducing rejections across both priority classes. This improvement comes at the cost of a modest increase in assignments to non-preferred physicians.
It is important to note that $\pi^{t}$ is computationally expensive. With 30 sampled future scenarios, the policy requires approximately 25 hours to complete 10 episodes when executed in parallel. $\pi^{*o}$ shows the best achievable performance with significantly low priority 1 patient rejections.
Policy Cost Priority 1 Rejected Priority 2 Rejected Undesirable Assignments
$\pi_{greedy}$ 3047.48 10.26 15.26 22.39
$\pi^{t}$ 1785.58 2.86 19.98 27.42
$\pi^{*o}$ 1271.25 0.45 20.45 25.32
Table 3: Average performance metrics for $\pi_{greedy}$, $\pi^{t}$ and $\pi^{*o}$
7.2.2 Learning without interaction (Isolated Learning)
We first report results obtained without expert interaction. The hyperparameter settings used for Algorithm˜1 are summarised in Table˜4 and correspond to the base configuration of the algorithm. Any deviations from this configuration are explicitly stated in the relevant experimental subsections. Setting $I=1$ implies that only a single outer iteration is performed, resulting in the training of a single learned policy.
We consider three expert policies: the full-information expert $\pi^{*o}$, the deterministic expert $\pi^{*p}$, and the two-stage stochastic expert $\pi^{*t}$. Each expert is queried once per state. To construct the dataset of expert demonstrations, we simulate a total of $H=4000$ episodes. Due to the imposed 30-second time limit on expert optimisation, only the two-stage stochastic expert $\pi^{*t}$ may terminate early $(\textbf{{MIP}}_{\textbf{SPPA}})$ and return a non-optimal action. As a result, the learned policy is ultimately trained using demonstrations generated by $\pi^{t}$. In contrast, both the full-information and deterministic experts can solve their respective optimisation models ($\textbf{{MIP}}_{\textbf{PPA}}$) to optimality within the allotted time.
The use of a time limit is necessary to keep the overall training time tractable. Since only a single learning iteration is performed, expert actions are selected using the vanilla DAgger decision rule. We also note that employing the full-information expert $\pi^{*o}$ within Algorithm˜1 is equivalent to solving deterministic instances of the PPA using a mixed-integer programming solver. Finally, to accelerate data collection, 10 episodes are simulated in parallel.
Hyperparameters
STOP() $I=1$
$F_{j}(.)$ 1
$\kappa$ 30 seconds
Decision Rule Vanilla DAgger
$\lambda$ 1
$\beta$ 0.8
Table 4: Hyper Parameters for Algorithm˜1
Full-Information and Deterministic Experts
We set $F_{e}(\cdot)$ to $\pi^{*o}$ and $\pi^{*p}$ for the full-information and deterministic expert configurations, respectively. Using Algorithm˜1, we collect approximately 400,000 expert demonstrations. Solving the $\textbf{{MIP}}_{\textbf{PPA}}$ requires, on average, 0.2 seconds per instance.
We first analyse the effect of the number of simulated episodes on the performance of the learned policy, to understand how the size of the demonstration dataset influences learning quality. The results are shown in Figure˜4. For the full-information expert, increasing the number of demonstrations does not lead to a clear or consistent improvement in performance. In contrast, policies trained using demonstrations from the deterministic expert exhibit a steady performance improvement as the number of demonstrations increases, ultimately achieving superior results.
A plausible explanation is that the deterministic expert encounters a wider range of system states, as it does not select actions that are optimal with respect to the realised trajectory. Consequently, the learned model is exposed to a more diverse state distribution, including states that would not typically be visited under an optimal policy. This increased diversity improves the generalisation capability of the learned policy when deployed. The black line in Figure˜4 represents the performance of the baseline policy $\pi^{t}$.
Figure 4: Test performance of learned models with different numbers of episodes NN
Table˜5 reports detailed results for the best-performing policies learned using the optimal full-information and deterministic experts. The higher cost relative to $\pi^{t}$ is primarily attributable to an increased number of rejected priority 1 patients under the policy learned from the full-information expert. The table also reports the total training time required to simulate 4,000 episodes.
It is important to note that the full-information expert can label all states within an episode in a single optimisation run by solving the $\textbf{{MIP}}_{\textbf{PPA}}$ once, and therefore does not require the use of Algorithm˜1 for data generation. In contrast, the deterministic expert must generate and solve a new optimisation instance at every state within an episode, which substantially increases the overall data collection time.
Learnt Policy Cost
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Full-Information ($\pi^{*o}$) 2031.34 3.63 22.19 25.77 0.5
Deterministic ($\pi^{*p}$) 1975.72 3.25 22.33 25.50 3.25
Table 5: Average performance metrics for Full-Information and Deterministic experts
Two-Stage Stochastic Expert
To employ the two-stage stochastic expert, with $F_{e}(\cdot)$ set to $\pi^{t}$, it is necessary to specify the number of future scenarios generated at each decision epoch. In this study, we use a fixed number of scenarios across all epochs. Alternative designs could allow the number of scenarios to vary over time—for example, by generating more scenarios in later stages, where the two-stage problem may be easier to solve and decisions are likely to be more critical with respect to the utilisation of remaining capacity.
We evaluate three configurations using 5, 15, and 30 scenarios. Figure˜5 illustrates the effect of the number of simulated episodes on the quality of the learned policy. The results indicate substantial performance improvements as the size of the training dataset increases. Performance gains are expected to plateau once the number of episodes exceeds approximately 2,000.
Figure 5: Test performance of learned models with different numbers of episodes NN
Table˜6 reports the results, showing that learning from a two-stage stochastic expert with multiple scenarios yields substantially better performance than learning from a full-information expert based on a single scenario. In particular, the learned policies significantly reduce the number of rejected patients. Among the configurations considered, the model trained with 30 scenarios performs best and achieves performance closest to that of the baseline policy $\pi^{t}$, as expected. These gains, however, come at the expense of increased training time, since multiple scenarios must be generated and a multi-scenario optimisation problem $(\textbf{{MIP}}_{\textbf{SPPA}})$ must be solved.
We further report detailed statistics on optimality gaps and solution times under the imposed 30-second time limit for expert calls in Table˜7. The table summarises results across approximately 400,000 expert calls. The majority of instances are solved to optimality or near optimality, as reflected by the 99% values. As the number of scenarios increases, a modest rise in the average optimality gap is observed. In cases where the optimality gap is non-zero, it is typically because the expert has identified the optimal solution but requires additional time to certify its optimality. Overall, the observed gaps suggest that the stochastic formulation remains relatively tractable and is capable of providing high-quality expert demonstrations.
With respect to solution times, the average solve time increases with the number of scenarios. Consequently, deploying $\pi^{t}$ with 30 scenarios for real-time decision-making would require approximately 5.09 seconds per decision on average. By contrast, the learned policy produces decisions within milliseconds. Solution times also vary across the planning horizon. Figure˜6 reports the average optimality gap and solution time by decision epoch. Relative to the overall averages, early decision epochs exhibit larger gaps and longer solution times, reflecting the greater problem size at the start of an episode. When using $\pi^{t}$ with a 30-second time limit, solution times at the beginning of an episode can approach 20 seconds on average. This highlights the importance of a fast reliable technique even more.
Learnt Policy Scenarios Average Cost
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Two-Stage ($\pi^{t}$) 5 1890.35 3.36 19.89 27.43 8.65
15 1822.45 3.03 20.02 27.38 32.47
30 1801.85 2.92 20.15 27.20 65.94
Table 6: Average performance metrics for Two-Stage Stochastic Expert
Scenarios Average Gap (%) 99% Gap (%) Average Solve Time(secs) 99% Solve Time(secs)
5 0 0 0.66 6.41
15 0.004 0.16 2.51 30
30 0.013 0.29 5.09 30
Table 7: Statistics across expert ($\pi^{t}$) calls
Figure 6: Average Gap and Average solve time across decision epoch
Aggregated Deterministic Expert
For a fixed number of scenarios, the deterministic expert, with $F_{e}(\cdot)$ set to $\pi^{*o}$, can be used to solve each scenario instance independently, after which the resulting solutions are aggregated to produce a single target action. Instance generation follows the same procedure as for the two-stage stochastic expert. Under this strategy, aggregation effectively serves as a heuristic approximation of a two-stage stochastic problem. Relative to the base configuration of Algorithm˜1, the primary modification is that the number of expert actions returned by $F_{j}(\cdot)$ is equal to the number of scenarios considered.
Figure˜7 illustrates the impact of the number of simulated episodes on the performance of the learned policy. Performance improves as the number of episodes increases during the initial phase of training. However, beyond approximately 500 episodes, additional data provide limited benefit, and in some cases, a marginal deterioration in performance is observed.
Figure 7: Test performance of learned models with different numbers of episodes NN
Table˜8 reports results across all scenario configurations considered. For aggregation, we convert the optimal actions into frequency vectors and use these as training targets for the ML model. This yields a smoother, continuous target representation, as opposed to the binary action targets used with the other expert configurations. Across all scenario counts, performance in terms of total cost remains broadly similar, indicating that increasing the number of scenarios offers limited additional benefit. Compared with policies learned from the two-stage stochastic expert, the aggregated deterministic policies perform worse overall.
Moreover, training time is substantially higher for the aggregated deterministic expert. This is because the expert must solve multiple optimisation problems sequentially to generate each target action, rather than solving a single multi-scenario formulation. As a result, the total training time exceeds that of the two-stage stochastic expert by a significant margin.
Learnt Policy Scenarios Average Cost
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Deterministic ($\pi^{*p}$) 5 1944.87 3.26 21.73 25.42 15.90
15 1950.09 3.28 21.72 25.55 50.63
30 1958.12 3.30 21.81 25.60 97.18
Table 8: Average performance metrics for Aggregated Deterministic Expert
7.2.3 Learning with interaction
We next allow interaction with the expert in order to iteratively learn decision policies. This requires modifying the stopping criterion of the learning procedure. Specifically, we impose a stopping criterion of $I=400$ learning iterations. In each iteration $i$, the current learned policy $\pi_{i}$ is executed for $H=10$ episodes. This results in a total of 4,000 simulated episodes used to construct the training datasets, matching the total number of episodes used in the isolated learning setting.
Since the single-scenario experts solve their underlying optimisation problems to optimality relatively quickly, they are able to support a large number of learning iterations. Upon completion of the learning process, we simulate all learned policies and select the policy that achieves the lowest average cost on the test instances.
Full-Information and Deterministic Experts
Figure˜8 reports the average cost across learning iterations for both expert types. The horizontal axis corresponds to the learned policy $\pi_{i}$ obtained after iteration $i$. The results show a consistent decrease in average cost as additional data are incorporated through expert interaction. Performance metrics for the best learned policies are summarised in Table˜9, with the corresponding values indicated by dotted lines in the figure.
The models learned through interaction achieve lower average costs compared with those obtained via isolated learning, as reported in Table˜5. Table˜9 also reports the percentage improvement relative to the isolated learning models. In addition, policies learned with interaction reject fewer patients overall. No significant performance difference is observed between policies learned from the full-information expert and those learned from the deterministic expert.
As shown in Figure˜8, learning performance begins to saturate after approximately 200 iterations, although the best-performing models are obtained at later iterations. We would like to point out that the best models obtained through isolated learning required 4000 training episodes. Using interaction, the learned model improves on those values within 2000 episodes. This highlights the effect of interaction.
Training times are also reported in Table˜9. Under interactive learning, the training time associated with the deterministic expert more than doubles. For the full-information expert, training time increases substantially due to the use of the vanilla DAgger decision rule. Under this rule, the states encountered during learning differ from those observed by a pure full-information expert, which necessitates solving the $\textbf{{MIP}}_{\textbf{PPA}}$ at each newly encountered state.
Figure 8: Test performance of different learned models
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Full-Information 1 261 1951.76 3.9 3.03 22.41 26.93 16.40
Deterministic 1 371 1943.85 1.6 3.14 22.25 25.57 14.77
Table 9: Average performance metrics for the best model using Full-Information and Deterministic Expert with interaction
Two-Stage Stochastic Expert
Figure˜9 reports the test performance of the learned policies across learning iterations. Compared with the full-information and deterministic expert settings, a stronger and more consistent decreasing trend in total cost is observed as learning progresses.
Table˜10 summarises the average performance of the best-performing policies obtained for each scenario configuration. These policies consistently outperform their isolated learning counterparts reported in Table˜6, as indicated by the dotted lines. Among all configurations, the policy trained using 30 scenarios achieves the best performance. It yields a marginal improvement over the policy learned without interaction, while requiring a significantly smaller number of expert demonstrations.
Relative to the baseline policy $\pi^{t}$, the learned policy trained with 30 scenarios achieves a comparable average cost. Moreover, it rejects fewer priority 1 patients. This improvement is accompanied by a slightly higher number of priority 2 rejections and an increased number of assignments to non-preferred physicians.
Figure 9: Test performance of different learned models
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Two-stage Stochastic 5 389 1844.75 2.4 2.98 20.56 27.21 23.19
15 301 1804.59 0.9 2.84 20.45 27.01 59.22
30 251 1795.93 0.3 2.81 20.41 27.13 101.83
Table 10: Average performance metrics for the best model using the Two-stage Stochastic Expert with interaction
Aggregated Deterministic Expert
Figure˜10 illustrates the performance of learned policies obtained using different numbers of aggregated scenarios, while Table˜11 reports the performance of the best-performing policies. No significant performance differences are observed across the scenario counts considered. Notably, the best-performing policy is obtained using aggregation over only five scenarios. This observation is consistent with the spread across scenario configurations shown in Figure˜10. Furthermore, the learned models perform marginally better than those obtained through isolated learning, while requiring a smaller number of expert demonstrations.
Overall, the learned policies outperform those trained using the full-information and deterministic experts. However, they do not achieve the same level of performance as policies learned from the two-stage stochastic expert. In addition, increasing the number of aggregated scenarios to 15 or 30 does not yield meaningful performance improvements relative to the isolated learning results reported in Table˜8.
Figure 10: Test performance of different learned models
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Aggregated Deterministic 5 366 1938.69 0.3 3.11 22.13 25.34 29.16
15 359 1941.93 0.4 3.17 21.98 25.53 64.42
30 190 1939.165 0.9 3.17 21.96 25.33 116.45
Table 11: Average performance metrics for the best model using the Aggregated Deterministic Expert with interaction
7.3 Inference
The experimental study compares different expert configurations in terms of interaction and the computational effort required to learn high-quality policies. Among all configurations, the policy learned from the two-stage stochastic expert with 30 scenarios achieves the best performance, both with and without interaction. Allowing expert interaction further improves policy quality, measured in terms of cost, when the same number of training episodes (4,000) is used. A clear learning trend is observed across iterations: when the initial dataset is empty, average costs are high in the early iterations, but decrease rapidly within the first 10 iterations. Beyond this initial phase, steady improvements are observed for all expert configurations as additional data are incorporated. Performance across consecutive iterations does not decrease monotonically, as the incremental data collected at each iteration induces only marginal updates to the weights of the multilayer perceptron.
From a computational standpoint, the full-information and deterministic experts enable the collection of a large volume of lower-quality demonstrations within relatively short computation times. For the PPA, the two-stage stochastic formulation remains tractable, allowing this expert to be used effectively by selecting an appropriate number of scenarios. However, as the number of scenarios increases, the time required to compute an expert action also increases. Imposing a strict time limit may therefore compromise the quality of the suggested action, particularly for more complex scheduling and routing problems.
The aggregated deterministic expert offers an alternative to the two-stage stochastic expert. For the PPA, however, policies learned using the aggregated deterministic expert do not outperform those learned from the two-stage stochastic expert. Moreover, aggregation requires repeated expert calls, which substantially increases the overall training time. Empirical results indicate that Algorithm˜1 with an aggregated deterministic expert, using the same number of scenarios, incurs a higher computational cost than the two-stage stochastic expert for the PPA. Ultimately, the choice between these approaches depends on the characteristics of the problem under consideration. As demonstrated in [gawas2024ildagger], the aggregated deterministic expert can outperform the two-stage stochastic expert in settings where the two-stage problem is more difficult to solve and high-quality solutions cannot be obtained within the imposed time limits. In such cases, the aggregated deterministic expert represents a viable and effective alternative.
8 Conclusion
This paper examined the role of the expert in imitation learning for sequential decision problems under uncertainty. While imitation learning has become a practical alternative to reinforcement learning in combinatorial optimisation, the notion of an expert has remained loosely defined in the literature. To address this gap, we proposed a structured taxonomy that characterises experts along key dimensions, including uncertainty modelling, optimality, interaction strategy, labelling approach, and trajectory control. This taxonomy provides a unified perspective for interpreting prior work and for guiding expert design in future studies.
Building on this framework, we introduced a DAgger-based algorithm that explicitly incorporates expert design choices and enables systematic comparison across expert configurations. An extensive experimental study on the physician–patient assignment problem illustrates how expert properties affect learning performance, data efficiency, and computational cost. The results reveal clear trade-offs between expert quality and scalability. Full-information and deterministic experts allow rapid collection of demonstrations but often yield lower-quality policies, whereas two-stage stochastic experts provide stronger supervision at the cost of increased computation.
The experiments further show that iterative interaction with the expert can improve the policy quality, particularly during early learning iterations. Although performance improvements are not strictly monotonic, the overall trend confirms the benefit of interaction in mitigating distribution shift. Among all configurations, policies learned from a two-stage stochastic expert with a sufficiently large number of scenarios achieve the strongest performance, closely matching optimisation-based baselines while enabling significantly faster decision making.
The study also evaluated an aggregated deterministic expert as a heuristic alternative to stochastic optimisation. While this approach can be useful when two-stage problems are difficult to solve within time limits, the results indicate that it does not outperform direct stochastic supervision for the considered application and leads to higher training costs due to repeated expert queries. These findings emphasise that expert selection should be guided by problem structure and computational constraints.
Overall, this work clarifies the importance of expert design in imitation learning for sequential decision making. By formalising expert characteristics and empirically analysing their impact, the paper provides both conceptual guidance and practical insights for deploying learning-based decision policies in complex operational settings. Future work may explore adaptive expert-query strategies, richer stochastic expert models, and applications to larger-scale optimisation problems.
Appendix
Appendix A MIP formulations
We first provide an MIP formulation for PPA for the static problem as well as the stochastic problem with $\Omega$ scenarios.
A.1 Deterministic MIP formulation
We give an MIP formulation for the PPA below. We denote the patients by $k$ and the physicians by $p$.
Table 12: Model, Parameters, and Variables
Sets
$\mathcal{P}$ set of all physicians
$\mathcal{K}$ set of all patients
$\mathcal{P}_{k}$ set of all eligible physicians for patient $k$
$\{1,2\}$ priority classes
Parameters
$P$ number of physicians
$K$ number of patients
$t_{k}$ duration of patient $k$
$r_{k}$ priority of patient $k$
$p_{k}$ preferred physician of patient $k$
$T$ total work capacity for physician $p$
$L_{p}$ total number of appointments allowed for physician $p$
$c_{r_{k}}^{\text{rej}}$ rejection cost for patient $k$
$c_{r_{k}}^{\text{pref}}$ penalty for assigning an undesired physician to patient $k$
Variables
$a_{kp}\in\{1,0\}$ $k\in\mathcal{K}$, $p\in\mathcal{P}$ 1 if patient $k$ is assigned to physician $p$
$u_{k}\in\{1,0\}$ $k\in\mathcal{K}$ 1 if patient $k$ is rejected
$\textbf{{MIP}}_{\textbf{PPA}}:=\min\quad\underbrace{\sum_{k\in\mathcal{K}}c^{\text{rej}}_{r_{k}}u_{k}}_{\text{Rejection cost}}\;+\;\underbrace{\sum_{k\in\mathcal{K}}\sum_{p\in\mathcal{P}}c^{\text{pref}}_{r_{k}}a_{kp}}_{\text{Non-preferred assignment penalty}}$ (16) s.t. $\sum_{p\in\mathcal{P}_{k}}a_{kp}+u_{k}=1$ $\forall k\in\mathcal{K}$ (17) $\sum_{k\in\mathcal{K}}a_{kp}\leq L_{p}$ $\forall p\in\mathcal{P}$ (18) $\sum_{k\in\mathcal{K}:p\in\mathcal{P}_{k}}t_{k}\,a_{kp}\leq T$ $\forall p\in\mathcal{P}$ (19) $a_{kp}\in\{0,1\}$ $\forall k\in\mathcal{K},\ \forall p\in\mathcal{P}$ (20) $u_{k}\in\{0,1\}$ $\forall k\in\mathcal{K}$ (21)
The objective function (16) is the sum of the total rejection cost and the penalty for all non-preferred assignments. Constraint (17) ensures that all patients are either assigned to a physician or rejected. Constraints (18) and (19) enforce capacity limits on the number of appointments and total workload, respectively.
A.2 Two-stage Stochastic MIP formulation
Table˜13 introduces additional notation for the two-stage formulation, where variables and parameters are augmented with a scenario index $\omega$. All other variables and sets retain the same meaning as in Table˜12. We use $\omega$ to denote a scenario. In a two-stage formulation, the first patient is common across all scenarios and is denoted by $k=1$.
Table 13: Model, Parameters, and Variables
Sets
$\Omega$ set of all scenarios
$\mathcal{K}_{\omega}$ set of all patients in scenario $\omega$
$\mathcal{P}_{k\omega}$ set of all eligible physicians for patient $k$ in scenario $\omega$
Parameters
$K_{\omega}$ number of patients in scenario $\omega$
$t_{k\omega}$ duration of patient $k$ in scenario $\omega$
$r_{k\omega}$ priority of patient $k$ in scenario $\omega$
$p_{k\omega}$ preferred physician of patient $k$ in scenario $\omega$
$c_{r_{k\omega}}^{\text{rej}}$ rejection cost for patient $k$ in scenario $\omega$
$c_{r_{k\omega}}^{\text{pref}}$ penalty for assigning an undesired physician to patient $k$ in scenario $\omega$
Variables
$a_{kp\omega}\in\{1,0\}$ 1 if patient $k$ is assigned to physician $p$ in scenario $\omega$
$u_{k\omega}\in\{1,0\}$ 1 if patient $k$ is rejected in scenario $\omega$
$\textbf{{MIP}}_{\textbf{SPPA}}:=\min\ \frac{1}{|\Omega|}\left[\sum_{\omega\in\Omega}\sum_{k\in\mathcal{K}_{\omega}}c^{\text{rej}}_{r_{k\omega}}u_{k\omega}+\sum_{\omega\in\Omega}\sum_{k\in\mathcal{K}_{\omega}}\sum_{p\in\mathcal{P}_{k\omega}}c^{\text{pref}}_{r_{k\omega}}a_{kp\omega}\right]$ (22) s.t. $\sum_{p\in\mathcal{P}_{k\omega}}a_{kp\omega}+u_{k\omega}=1$ $\forall\omega\in\Omega,\ \forall k\in\mathcal{K}_{\omega}$ (23) $\sum_{k\in\mathcal{K}_{\omega}:p\in\mathcal{P}_{k\omega}}a_{kp\omega}\leq L_{p}$ $\forall\omega\in\Omega,\ \forall p\in\mathcal{P}$ (24) $\sum_{k\in\mathcal{K}_{\omega}:p\in\mathcal{P}_{k\omega}}t_{k\omega}\,a_{kp\omega}\leq T$ $\forall\omega\in\Omega,\ \forall p\in\mathcal{P}$ (25) $a_{1p\omega}=a_{1p\omega^{\prime}}$ $\forall\omega\neq\omega^{\prime}$ (26) $u_{1\omega}=u_{1\omega^{\prime}}$ $\forall\omega\neq\omega^{\prime}$ (27) $a_{kp\omega}\in\{0,1\}$ $\forall\omega\in\Omega,\ \forall k\in\mathcal{K}_{\omega},\ \forall p\in\mathcal{P}$ (28) $u_{k\omega}\in\{0,1\}$ $\forall\omega\in\Omega,\ \forall k\in\mathcal{K}_{\omega}$ (29)
The objective (22) and constraints (23)–(25) mirror those of $\textbf{{MIP}}_{\textbf{PPA}}$. Constraints (26) and (27) enforce non-anticipativity of the first-stage decisions across scenarios.
Appendix B Dirichlet-based eligibility construction
Eligible physicians are generated using a patient-level preference model based on the Dirichlet distribution. Let $\mathcal{P}=\{1,\dots,P\}$ denote the physician set. For each arriving patient, a probability vector
$\mathbf{w}=(w_{1},\dots,w_{P}),\qquad w_{p}\geq 0,\ \sum_{p\in\mathcal{P}}w_{p}=1,$
is specified as a global baseline over physicians. In the experiments with $P=4$, we use $\mathbf{w}=[0.4,0.3,0.15,0.15]$. A concentration parameter $\alpha>0$ controls how strongly patient preferences concentrate around $\mathbf{w}$. We form the Dirichlet parameter vector
$\boldsymbol{\alpha}=\alpha\,\mathbf{w},$
and sample a patient-specific taste vector
$\mathbf{z}=(z_{1},\dots,z_{P})\sim\mathrm{Dirichlet}(\boldsymbol{\alpha}).$
Each component $z_{p}$ represents the relative affinity of the current patient for physician $p$, with $\sum_{p}z_{p}=1$. Larger values of $\alpha$ produce tastes that are closer to the baseline $\mathbf{w}$, whereas smaller values yield more dispersed and heterogeneous tastes across patients. In our implementation, $\alpha=25$ is used when the baseline weights are enabled, and $\alpha=1$ is used when all physicians are treated uniformly (that is, $w_{p}=1/P$ for all $p$).
Given $\mathbf{z}$, the eligible set size is drawn as
$K\sim\mathcal{U}\{k_{\min},\dots,k_{\max}\},$
and the eligible physician set for the patient is defined by selecting the $K$ physicians with the largest taste values:
$\mathcal{P}_{k}=\operatorname{TopK}(\mathbf{z},K).$
Equivalently, we construct an eligibility indicator $E_{p}\in\{0,1\}$ for each physician $p$, where $E_{p}=1$ if and only if $p\in\mathcal{P}_{k}$. A safety check enforces that at least one physician is eligible, which is guaranteed by construction for $K\geq 1$.
After constructing $\mathcal{P}_{k}$, the preferred physician is selected from within the eligible set using the taste vector restricted to $\mathcal{P}_{k}$. Specifically, the preferred physician is sampled with probabilities proportional to $\{z_{p}:p\in\mathcal{P}_{k}\}$, so that physicians with higher taste are more likely to be preferred while maintaining feasibility with respect to eligibility.