### v2

# Imitation Learning for Combinatorial Optimisation under Uncertainty

Abstract
Imitation learning (IL) provides a data-driven framework for approximating policies for large-scale combinatorial optimisation problems formulated as sequential decision problems (SDPs), where exact solution methods are computationally intractable. A central but underexplored aspect of IL in this context is the role of the expert that generates training demonstrations. Existing studies employ a wide range of expert constructions, yet lack a unifying framework to characterise their modelling assumptions, computational properties, and impact on learning performance.
This paper introduces a systematic taxonomy of experts for IL in combinatorial optimisation under uncertainty. Experts are classified along three dimensions: (i) their treatment of uncertainty, including myopic, deterministic, full-information, two-stage stochastic, and multi-stage stochastic formulations; (ii) their level of optimality, distinguishing task-optimal and approximate experts; and (iii) their interaction mode with the learner, ranging from one-shot supervision to iterative, interactive schemes. Building on this taxonomy, we propose a generalised Dataset Aggregation (DAgger) algorithm that supports multiple expert queries, expert aggregation, and flexible interaction strategies.
The proposed framework is evaluated on a dynamic physician-to-patient assignment problem with stochastic arrivals and capacity constraints. Computational experiments compare learning outcomes across expert types and interaction regimes. The results show that policies learned from stochastic experts consistently outperform those learned from deterministic or full-information experts, while interactive learning improves solution quality using fewer expert demonstrations. Aggregated deterministic experts provide an effective alternative when stochastic optimisation becomes computationally challenging.
keywords:
Imitation Learning, Operations Research, Sequential Decision Making
category:
[…]
1 Introduction
Combinatorial optimisation (CO) lies at the core of many real-world decision-making problems, including scheduling, logistics, network design, and resource allocation. These problems are challenging due to the combinatorial explosion in the number of feasible solutions, whereby even relatively small instances can generate an enormous search space. This difficulty is further compounded by complex constraints and the requirement to identify globally optimal solutions rather than merely feasible ones. As problem size increases, naïve or brute-force approaches quickly become impractical, necessitating more sophisticated methods for efficiently exploring the solution space.
The computational burden becomes particularly severe in settings involving uncertainty or temporal dynamics, such as stochastic or sequential decision problems, where multiple future scenarios must be considered. In such cases, exact optimisation methods (such as branch-and-bound or dynamic programming) often become computationally intractable. This challenge is further exacerbated in applications that require fast and reliable decision-making. To address these limitations, practitioners rely on a variety of approximation techniques, including heuristics, metaheuristics, and problem-specific algorithms. These approaches trade off optimality guarantees for computational efficiency, thereby enabling the practical deployment of CO methods in complex operational environments.
Machine learning (ML) has recently emerged as a powerful complementary tool for tackling CO problems, particularly in regimes where traditional exact or heuristic methods scale poorly. In many operational settings, decisions must be made repeatedly and under strict time constraints, making it impractical to solve large optimisation models from scratch at every decision point. In such environments, ML methods can exploit historical data and structural regularities in problem instances to learn approximate decision rules or policies that generalise across scenarios. To illustrate this motivation, consider problems such as assigning incoming patients to physicians, dispatching vehicles to delivery requests, or selecting branching decisions within a mixed-integer solver. Solving a full optimisation model at every decision point can be computationally prohibitive, particularly in real-time applications. Yet, across repeated instances, consistent structural patterns often emerge in how high-quality decisions depend on the system state. Machine learning models can exploit these patterns by learning from historical solutions or expert demonstrations, allowing them to approximate the outcome of the optimisation process efficiently. Once trained, such models can generate near-optimal decisions in milliseconds rather than seconds or minutes, while remaining responsive to evolving system conditions.
Reinforcement learning (RL) [sutton1998reinforcement] has demonstrated strong potential, particularly in learning decision policies through interaction with an environment, utilising reward feedback to improve performance iteratively. Such approaches are well-suited to settings where explicit models are unavailable, incomplete, or prohibitively complex. Alongside RL, Imitation learning (IL) [hussein2017imitation] has gained attention as an attractive alternative, especially in scenarios where defining a suitable reward function is difficult but expert demonstrations are available. Instead of learning through trial and error, IL aims to learn a direct mapping from system states to actions by mimicking expert behaviour. A central challenge in this paradigm is the acquisition of high-quality expert demonstrations, which are often expensive or infeasible to generate due to the computational cost of solving CO problems optimally. This has motivated growing interest in approximate, surrogate, or interactive expert models that can provide informative guidance while remaining computationally tractable. The design and utilisation of such experts remain key open questions at the intersection of ML and combinatorial optimisation.
Contribution
The idea of imitating expert policies in CO is relatively recent, and existing work employs a wide range of strategies to define or approximate experts. These include using offline optimal solutions, simplified surrogate models, and structured or repeated interactions with problem-specific oracles. We first survey recent literature to illustrate the breadth of approaches to expert construction and utilisation. However, despite this growing body of work, there is limited understanding of what constitutes an expert in IL for CO and how different expert design choices influence learning outcomes.
To address this gap, we introduce a taxonomy of expert design for IL in CO. The proposed taxonomy unifies existing approaches by organising experts along key conceptual dimensions, providing a coherent framework that connects and contextualises a wide range of prior methods. It also serves as a guide for the principled selection and design of experts across different problem settings.
In the second part of the paper, we propose a general two-stage learning pipeline in which a CO-based expert layer generates demonstrations that are then used to train an ML-based decision model. The framework exposes several design choices that correspond directly to the dimensions of the expert taxonomy. We conduct a comparative empirical study using a physician scheduling problem to examine how these choices affect learning performance. The results offer practical insights into the trade-offs between expert quality, computational effort, and solution quality in sequential decision-making problems under uncertainty.
Scope
Several existing surveys examine the integration of ML and CO [bengio2021machine, kotary2021end, mandi2024decision, sadana2025survey, scavuzzo2024machine]. In contrast, this article focuses specifically on the role and design of the expert in IL. We concentrate on settings involving sequential decision making under uncertainty, where expert queries are costly and real-time decisions are required. Nevertheless, we also consider studies involving static and two-stage stochastic problems for completeness. Experts may take the form of human decision makers, problem-specific heuristics, or exact optimisation solvers. The overarching objective is to develop learning-based models that approximate or replace expert decision making, thereby enabling scalable and timely decisions without substantial loss in solution quality. No assumptions are made regarding specific structural properties of the underlying optimisation problems.
Organisation
The remainder of the paper is organised as follows. Section 2 reviews key concepts in ML and CO for sequential decision making. Section 3 introduces the fundamentals of IL. Section 4 presents the proposed taxonomy of expert policies and situates it within the existing literature. Section 6 describes the problem setting considered in this study, and Section 5 details the proposed methodology for applying IL to sequential decision making under uncertainty. Section 6.2 presents the experimental results and analysis. Section 8 concludes the paper.
2 Background
Many real-world CO problems can be naturally formulated as sequential decision problems (SDPs) [powell2007approximate]. At each decision epoch, an agent selects an action based on the current system state and available information, while accounting for both immediate consequences and future system evolution. The process is inherently dynamic, as new information is revealed and the system transitions over time. The objective is to design policies that minimise cumulative cost or maximise cumulative reward over a finite or infinite planning horizon, making SDPs well-suited to problems involving uncertainty, timing, and adaptive decision-making.
2.1 Markov Decision Process
SDPs are commonly modelled using the Markov decision process (MDP) framework, which provides a formal representation of decision-making under uncertainty [powell2022reinforcement]. An MDP models the interaction between controllable actions and stochastic system dynamics over time. In Table˜1, we summarise the core components of an MDP. Additional details are provided in the appendix.
Table 1: Key components of a Markov decision process
Component Description
Decision epoch ($k$) A time index at which the decision maker observes the system state and selects an action. We focus on finite-horizon problems with final epoch $K<\infty$.
State ($x_{k}$) A complete description of the system at epoch $k$ that is sufficient for future decision making. The set of all states defines the state space $\mathcal{S}$.
Action ($a_{k}$) The decision taken at epoch $k$, which may be discrete, continuous, or vector-valued, and must satisfy feasibility constraints $a_{k}\in\mathcal{A}_{k}$.
Exogenous information ($\xi_{k}$) Random information revealed after action selection, representing uncertainty external to the system.
Cost/Reward function ($C_{k}(x_{k},a_{k})$) The immediate cost incurred or reward received when an action is taken in a given state at epoch $k$.
Transition function ($X^{M}(x_{k},a_{k},\xi_{k})$) The mapping that governs the evolution of the system from one state to the next, capturing the system dynamics.
Policy ($\pi$) A mapping from states to admissible actions. Policies may be deterministic or parameterised functions.
A particularly important special case arises when $K=1$. When uncertainty is present, this setting corresponds to a two-stage stochastic programming problem with a single realisation of uncertainty. In the absence of uncertainty, it reduces to a deterministic optimisation problem. Although such problems are static in nature, they often arise repeatedly in practice. For example, a company may solve an independent delivery routing problem each day, with no direct dependence across days. When $K>1$, the problem becomes a multi-stage or fully SDP.
2.2 Machine Learning
Machine learning (ML) [hastie2009elements, bishop2006pattern] is a subfield of artificial intelligence concerned with learning patterns from data to make predictions or decisions without explicit programming. At a high level, the objective is to learn a mapping from inputs to outputs. Let $X$ denote the input (feature) space and $Y$ the output space. We consider parameterised models of the form $f_{\theta}:X\rightarrow Y$, where $\theta\in\Theta$ denotes the model parameters. Learning consists of selecting parameters that yield an effective input–output mapping based on observed data and a chosen performance criterion. Depending on the type of feedback available, learning problems are commonly classified into supervised, unsupervised, and reinforcement learning.
Supervised learning considers settings in which the learner is given a finite collection of labelled examples $\{(x_{i},y_{i})\}_{i=1}^{N}$, with $x_{i}\in X$ and $y_{i}\in Y$. The objective is to approximate the underlying input–output relationship by minimising a loss function $\ell:Y\times Y\rightarrow\mathbb{R}$ that measures prediction error. A common example is the mean squared error,
$\frac{1}{N}\sum_{i=1}^{N}\left(f_{\theta}(x_{i})-y_{i}\right)^{2}.$
Since the true data-generating distribution is unknown, optimisation is performed empirically over the observed samples.
Unsupervised learning addresses scenarios in which only input samples $\{x_{i}\}_{i=1}^{N}$ are available, without labels. The goal is to uncover latent structure or informative representations of the data. Typical tasks include clustering, dimensionality reduction, and representation learning. Evaluation is generally based on internal metrics or qualitative assessment, as ground truth labels are unavailable.
Reinforcement learning focuses on sequential decision-making problems in which an agent interacts with an environment over time. At each step, the agent observes a state, selects an action, and receives scalar feedback in the form of a reward, together with a new state. Unlike supervised learning, explicit target outputs are not provided; instead, the agent must infer desirable behaviour from delayed and possibly sparse rewards. The objective is to learn a policy that maximises expected cumulative reward, typically formalised using the MDP framework.
The learning process involves several interconnected components. First, a model class is selected, such as linear predictors, decision trees, kernel methods, or neural networks, each offering different expressive capabilities. Learning proceeds by specifying an appropriate loss function and employing an optimisation algorithm, such as gradient descent or its variants, to adjust model parameters to minimise the loss on the training data. In addition to model parameters, learning algorithms involve hyperparameters, such as learning rates, regularisation strengths, or architectural choices, which govern the learning dynamics and model complexity. To evaluate generalisation performance and mitigate overfitting, data are typically partitioned into a training set, a validation set for hyperparameter tuning, and a test set for final evaluation. Overfitting occurs when a model fits the training data well but fails to generalise to unseen instances. We refer the reader to [hastie2009elements] for more information of these elements.
2.3 Machine Learning for MDPs
Machine learning provides several paradigms for addressing MDPs in settings where system dynamics are partially known or entirely unknown. Two prominent approaches are reinforcement learning (RL) and approximate dynamic programming (ADP). RL is typically applied in model-free settings, where transition dynamics are unknown, and policies are learned directly through interaction with the environment. Methods such as Q-learning and policy gradient algorithms rely on sampled trajectories to estimate value functions or policy gradients and to iteratively improve decision policies.
Approximate dynamic programming, which has its roots in operations research and control, often assumes access to a simulator or generative model of the system. ADP mitigates the intractability of exact dynamic programming by approximating value functions or policies using function approximation techniques, including linear basis functions, regression trees, and neural networks. These approximations enable iterative solution methods such as fitted value iteration and approximate policy iteration. Although RL and ADP differ in historical origins and terminology, the distinction between the two has become increasingly blurred, particularly with the adoption of deep learning–based methods.
IL provides an alternative paradigm for solving MDPs, especially in settings where specifying a reward function is difficult, misaligned with operational objectives, or undesirable from a modelling perspective. Rather than learning through trial-and-error interaction, IL seeks to learn a policy directly from expert-generated state–action trajectories. This expert-driven formulation is particularly appealing in combinatorial optimisation settings, where high-quality decision policies may be available through optimisation solvers or heuristics, but reward design and exploration are costly.
The effectiveness of IL depends critically on the quality, availability, and structure of the expert demonstrations. While simple approaches such as behavioural cloning treat imitation as a supervised learning problem, more advanced methods incorporate expert interaction or infer latent objectives. The design and use of experts, therefore, play a central role in applying IL to MDPs arising from complex optimisation problems.
The focus of this work is the application of IL to CO problems under uncertainty, with particular emphasis on the role and design of the expert. Building on the formulation of MDPs and the use of expert demonstrations, the following section provides a detailed overview of IL techniques.
3 Imitation Learning: Models and Challenges
IL [hussein2017imitation] is a specialised paradigm within supervised ML in which an agent learns to perform a task by mimicking behaviour demonstrated by an expert. Unlike classical supervised learning, where data points are typically assumed to be independent and identically distributed, IL commonly arises in sequential decision-making settings. In such problems, observations are temporally correlated and depend on previously executed actions.
The term expert refers to the source of demonstrations used by the learning agent and may correspond to a human decision maker, a heuristic, or an exact optimisation algorithm. Demonstrations are typically provided as trajectories $\tau=\{x_{i},a_{i}^{*}\}_{i=1}^{T}$, where $T$ denotes the trajectory length. An expert dataset is then defined as $\mathcal{D}=\{\tau_{j}\}_{j=1}^{N}$, with each state $x_{i}$ generated under the expert policy $\pi^{*}$ and the corresponding action given by $a_{i}^{*}=\pi^{*}(x_{i})$. In an ML context, actions are also referred to as targets or labels. Standard supervised learning methods can be applied to $\mathcal{D}$ to train a model that predicts expert-like actions from observed states. Here, each state–action pair is an independent training example, effectively flattening trajectories, but it does not require explicitly uncoupling or destroying the underlying trajectory structure.
IL provides a data-driven approach for solving complex SDPs with combinatorial action spaces. However, its effectiveness depends critically on access to high-quality expert demonstrations, which are often expensive to obtain when expert actions are generated by optimisation solvers. In addition, IL must scale to high-dimensional decision spaces and address covariate shift, whereby the learned policy encounters states at deployment that differ from those observed during training, leading to compounding errors. This section introduces two core IL paradigms, discusses these challenges, and motivates the use of approximate or imperfect experts.
3.1 Learning Paradigms
We describe two principal learning paradigms within IL: behavioural cloning and inverse RL, followed by a discussion of practical challenges and common mitigation strategies.
3.1.1 Behavioural Cloning (BC)
Behavioural cloning [pomerleau1988alvinn] formulates imitation as a supervised learning problem that maps states or observations directly to expert actions. The objective is to learn the parameters $\theta$ of a policy $\pi_{\theta}$ by solving
$\theta=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{(x,a^{*})\in\mathcal{D}}\big[\mathcal{L}(x,a^{*},\pi_{\theta})\big].$ (1)
Equation (1) minimises the one-step deviation between the learner and the expert along demonstrated trajectories. In the context of CO, two behavioural cloning formulations are commonly employed.
  • Parameter Prediction Task (PPT) - In this approach, an ML model predicts parameters that define a CO problem, which is subsequently solved by an external optimiser. The idea is to predict parameters that produce solutions close to the true ones in an imitation learning setting, where the expert knows the true parameters. The training objective is typically expressed in terms of regret (also referred to as task loss in [donti2017task]) when the predicted parameters appear in the objective of the problem. Regret is defined as the difference between the objective value achieved by the predicted decision and that of the expert decision under a realisation of uncertainty $\xi$. Following [sadana2025survey], the regret loss is given by
    $\mathcal{L}_{R}(x,a^{*},\pi_{\theta})=C(x,\pi_{\theta}(x),\xi)-C(x,a^{*},\xi),$ (2)
    where $C$ denotes the cost function. The loss in (2) is non-negative and vanishes only when the predicted decision coincides with the expert action. This paradigm is also referred to as the regret minimisation task (RMT) [sadana2025survey]. When predicted parameters appear in the constraints of the problem, regret alone is not sufficient, and one also needs to minimise the likelihood of infeasible outcomes. In such cases, the loss function involves a combination of a penalty for having infeasibility and losing optimality [mandi2025feasibility].
  • Action Imitation Task (AIT) - In contrast, AIT trains a model $\pi_{\theta}$ to directly output an action that approximates the expert decision, thereby eliminating the need for an optimisation solver at inference time. A common imitation loss is
    $\mathcal{L}_{I}(x,a^{*},\pi_{\theta})=\ell(\pi_{\theta}(x),a^{*}),$ (3)
    where $a=\pi_{\theta}(x)$ and $\ell(\cdot,\cdot)$ may be mean squared error or binary cross-entropy. When regret is used as the distance metric, AIT closely aligns with PPT. This approach is also known as learning to optimise (LtO) [kotary2024learning].
Comparing PPT and AIT
Under the PPT paradigm, the ML model predicts parameters of an easier CO problem that is subsequently solved by an external optimiser. By contrast, AIT eliminates this two-stage inference pipeline by training the model to directly output decisions, avoiding solver calls at deployment time. Nevertheless, an optimisation solver remains necessary during data generation to compute expert actions $a^{*}$. As a result, both paradigms rely on expert optimisation, but differ in when optimisation is used.
Figure 1: PPT
[…]
Figure 2: AIT
$\boldsymbol{x}$$\min_{a\in\mathcal{F}}C(x,a,\bar{\xi})$$\pi_{\theta}$$\boldsymbol{\hat{a}}$  $\boldsymbol{\hat{a}}$Refer to captionRefer to captionOptimization PhasePrediction Phase
IL with AIT models in CO settings presents several challenges, primarily due to large and complex decision spaces and the presence of ambiguous demonstrations. Without explicitly structured output spaces, standard models may generate infeasible or inconsistent actions. Feasibility can be restored through post-processing techniques such as projection or heuristic correction, although this may result in a modest increase in regret [kotary2024learning].
The PPT approach can mitigate infeasibility when the predicted parameters appear only in the objective function. However, when predicted parameters enter the constraints, the resulting solution may violate feasibility with respect to the constraints defined by the true parameters. Similar to AIT, [hu2023predict+] generate an initial solution using predicted parameters and subsequently apply corrective actions, if necessary, at an additional cost.
A key challenge in PPT lies in differentiating the loss function, as it depends on the optimisation outcome. Common strategies to address this include regularisation, perturbation techniques, and the use of surrogate loss functions [mandi2024decision, sadana2025survey]. Both PPT and AIT therefore offer promising approaches for decision making in CO under uncertainty; however, a systematic understanding of their relative strengths and limitations remains limited.
Distribution Shift
- Given an offline training dataset $\mathcal{D}=\{(x_{i},a_{i}^{*})\}_{i=1}^{N}$ generated by an expert policy, that is $(x_{i},a_{i}^{*})\sim\pi^{*}$, behavioural cloning seeks to solve Equation˜1. This formulation gives rise to a distribution shift, or distribution mismatch, a phenomenon well documented [ross2011reduction] in the supervised learning literature for sequential decision-making problems. During deployment, deviations from expert behaviour may lead the learned policy to encounter states that were absent from the training data. Such deviations can compound over time, causing the policy to drift into unfamiliar regions of the state space where reliable recovery is difficult.
To achieve stronger generalisation, one would instead aim to solve
$\theta=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{x\sim\pi_{\theta}}\big[\mathcal{L}(x,a^{*},\pi_{\theta})\big],$ (4)
thereby training the model under the state distribution induced by its own actions. This formulation exposes the learner to states arising from suboptimal decisions and enables exploration of a broader portion of the transition dynamics, improving robustness at deployment.
A standard methodology for addressing this mismatch is dataset aggregation (DAgger) Algorithm˜1, introduced by [ross2011reduction]. DAgger iteratively generates trajectories using the current learned policy $\pi_{i}$ and labels the visited states with actions provided by the expert (steps 4 and 5). The action to be executed at any epoch is given by a function called the decision rule ($DR(\cdot)$) in step 7. An example of the Vanilla DR is given in Algorithm˜2. Vanilla DAgger executes the expert action with a decaying probability and the learner action otherwise, gradually shifting control from the expert to the learned policy over training iterations. At the end of the episode, a new policy denoted by $\hat{\pi}_{i+1}$ is trained on the updated dataset. By repeatedly collecting expert labels on states visited by the learner, DAgger aligns the training distribution with the state distribution induced at deployment. However, a key limitation in CO settings is the high computational cost of step 5, as expert labelling typically involves repeatedly solving CO problems at each decision epoch. We revisit the DAgger algorithm in Section˜5.
1
2Function DAgger():
3   Initial policy $\hat{\pi}_{0}$; expert $\pi^{*}$; iterations II; horizon $K$; learning routine $\textsc{Train}(\cdot)$; $\mathcal{D}\leftarrow\emptyset$ for $i\leftarrow 0$ to I1I-1 do
4      Observe initial state $x_{0}$;
    /* Collect a K-step trajectory */
5    for $k\leftarrow 0$ to K1K-1 do
       $a^{*}_{k}\leftarrow\pi^{*}(x_{k})$     // Query expert action
6
       $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x_{k},a^{*}_{k})\}$  // Append demonstration
7
       $a_{k}\leftarrow DR(x_{k},a^{*}_{k},\hat{\pi}_{i})$  // Select action using decision rule
8
       $x_{k+1}\leftarrow X^{M}(x_{k},a_{k},\xi_{k})$// Environment transition
9
10      end for
11
    /* Retrain policy on aggregated dataset */
12    $\hat{\pi}_{i+1}\leftarrow\textsc{Train}(\mathcal{D})$;
13
14   end for
15
16 end
Algorithm 1 DAgger (Dataset Aggregation)
1Function VanillaDAgger($x_{k},{a}_{k},\hat{\pi}_{i}$):
2
3 $\beta_{i}\leftarrow\lambda^{i-1}\beta_{0}$;
4 $z\sim\text{Uniform}(0,1)$;
5
6 if $z\leq\beta_{i}$ then
7    return ${a}_{k}$;
8
9   end if
10 return $\hat{\pi}_{i}(x_{k})$;
11
12 end
Algorithm 2 VANILLA DAgger Decision Rule
3.1.2 Inverse Reinforcement Learning (IRL)
IRL, first introduced by [russell1998learning], aims to recover the underlying reward function that explains expert behaviour, rather than directly learning a mapping from states to actions. Following standard practice in the IRL literature, we adopt reward-based notation, in contrast to the cost-based formulation used for behavioural cloning. Once the reward function is inferred, a policy can be obtained by solving the resulting RL problem.
Formally, IRL seeks a reward function parameterised by $\theta$ such that the optimal policy induced by this reward coincides with the expert policy $\pi^{*}$:
$\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}\Bigg[\sum_{t=1}^{T}r_{\theta}(x_{t},\pi(x_{t}))\Bigg].$ (5)
Given an expert dataset $\mathcal{D}$, the reward parameters are estimated by solving
$\theta=\operatorname*{arg\,max}_{\theta}\;\mathcal{J}(\mathcal{D},r_{\theta}).$ (6)
Because policy optimisation in IRL is typically performed using RL, IRL methods naturally account for long-term consequences of actions and are therefore less sensitive to covariate shift than behavioural cloning. However, this robustness comes at a high computational cost. IRL is generally sample inefficient, requires repeated interaction with the environment, and alternates between reward estimation and policy optimisation. Moreover, IRL is inherently ill-posed, since multiple reward functions may induce the same expert behaviour, making the unique recovery of the true reward function impossible in general.
A variety of IRL methods have been proposed, differing primarily in how the objective function $\mathcal{J}$ is defined and optimised. Notable approaches include maximum margin methods, which separate expert and non-expert trajectories; maximum entropy IRL, which matches expert behaviour while encouraging stochasticity; Bayesian IRL, which infers a distribution over reward functions; and generative adversarial imitation learning (GAIL), which frames imitation as an adversarial learning problem. Comprehensive surveys of these methods are provided in [zare2024survey] and [zheng2022imitation].
In practice, IRL is most suitable when expert demonstrations are limited and accurate modelling of long-term behaviour is essential. A key advantage is that once a reward function is learned, policies can be re-optimised under changing system dynamics without additional demonstrations. By contrast, behavioural cloning is preferable when expert data are abundant and sufficiently cover the relevant state–action space, and when small deviations from expert behaviour do not lead to severe performance degradation.
3.2 Learning from Noisy or Suboptimal Experts
Although IL relies on expert demonstrations, fully optimal and noise-free data are rarely available in practice. Human experts may be inconsistent, and computing optimal solutions can be prohibitively expensive. Consequently, most real-world IL applications rely on imperfect but readily available demonstrations. In CO, human experts are generally unsuitable due to problem complexity, and optimisation solvers are used instead. However, solving SDPs under uncertainty is computationally demanding [powell2022reinforcement], motivating the use of static approximations or simplified scenarios to generate demonstrations.
Several studies in classical IL literature adopt this approach by solving static or simplified CO problems offline to obtain training targets [pham2023prediction, baty2024combinatorial]. Although such solutions are suboptimal, they are informative and significantly cheaper to compute. In many cases, these demonstrations are treated as optimal and used directly within classical IL pipelines. Nevertheless, given the $\mathcal{NP}$-hard nature of many CO problems, even static instances often require heuristics, time limits, or prescribed optimality gaps, introducing additional noise into the demonstrations. Despite its prevalence, the impact of such noise on IL performance in CO settings remains poorly understood, and systematic comparative studies are limited.
Recent work has begun to address learning under noisy supervision explicitly. For example, [xu2022discriminator, kim2021demodice] propose offline IL frameworks that leverage supplementary datasets generated by suboptimal policies. [sekhari2024selective] develop an IL algorithm that reduces expert queries while achieving low regret under noisy demonstrations. [wang2023imitation] introduce a two-stage approach that denoises demonstrations using diffusion models before imitation. Additional contributions in this area include [wu2019imitation, sasaki2020behavioral, wang2021learning], as summarised in [zare2024survey].
An additional challenge arises when multiple labels are available for the same state, often generated by different imperfect experts. Such situations may lead to conflicting demonstrations—for instance, when different drivers adopt distinct strategies under identical conditions. While this diversity can enrich the learning signal, it necessitates aggregation or filtering mechanisms to avoid degraded policy performance. Early work by [raykar2009supervised, raykar2010learning] addresses this issue in pre-labelled datasets. More recently, [sun2023mega] propose MEGA-DAgger, an interactive framework for learning from multiple imperfect experts that filters unsafe demonstrations and evaluates experts using scenario-specific criteria. [sekhari2024selective] further extend this line of work to settings in which different experts specialise in distinct regions of the state space. However, this line of work has not been explored when using IL for problems in operations research.
3.3 Applications of IL in SDPs
Recent applications of ML to SDPs in CO broadly fall into two categories. The first, ML-augmented optimisation, uses learning models to enhance classical optimisation algorithms. The second, known as end-to-end learning (E2EL), integrates optimisation components directly within learning architectures. This distinction follows the taxonomy of [kotary2021end]. While both IL and RL are used in these settings, we focus on work that applies IL to guide decision making within these two paradigms.
3.3.1 ML-Augmented Optimisation
ML-augmented optimisation embeds learning models within optimisation algorithms to guide or accelerate key decisions. Learning may be used to initialise algorithms, tune parameters, or repeatedly inform decisions during execution, such as branching, node selection, or column generation [bengio2021machine]. Existing approaches either predict discrete decisions directly or estimate surrogate scores that rank candidate actions and guide the search process.
  1. 1.
    Branch-and-Bound - We refer the reader to [bengio2021machine] and [scavuzzo2024machine] for a basic introduction to Branch-and-Bound (B&B). Branch-and-Bound (B&B) relies on two core decisions: selecting a branching variable and choosing a node to expand. Several studies use IL to guide these decisions. Works such as [alvarez2017machine, alvarez2016online, balcan2018learning, ding2020accelerating, gasse2019exact, gupta2022lookback, khalil2016learning, lin2022learning, lu-2020, Nair2020SolvingMI, zarpellon2021parameterizing] apply IL to learn branching policies. In parallel, [he2014learning, khalil2022mip, labassi2022learning, yilmaz2021study] focus on node selection strategies. The DAgger algorithm is employed in [he2014learning] and [Nair2020SolvingMI] to incorporate expert interaction. Additional contributions include learning when to invoke primal heuristics [khalil2017learning] and predicting backdoors to accelerate mixed-integer programming [cai2024multi].
  2. 2.
    Cutting Planes - Cut generation is a central technique in mixed-integer programming, aimed at tightening relaxations without excluding feasible integer solutions. Several studies use supervised learning to decide which cuts to add, including [deza2023machine, ding2020accelerating, huang2022learning, paulus2022learning].
  3. 3.
    Column Generation and Branch-and-Price - Column generation solves large-scale linear programmes by generating variables on demand. When combined with Branch-and-Bound, this yields the Branch-and-Price framework [barnhart1998branch]. Learning-based methods in this setting focus on selecting promising columns or designing effective branching strategies to improve speed. Representative works include [furian2021machine, morabit2021machine, pereira2022learning, sun2022learning, shen2022enhancing].
  4. 4.
    Large Neighbourhood Search - Large Neighbourhood Search (LNS) improves solutions by iteratively destroying and repairing partial solutions. ML has been used to generate high-quality initial solutions and to adaptively select neighbourhoods [hizir2025large, huang2023searching, liu2022learning, mossina2019multi, Nair2020SolvingMI, song2020general, sonnerat2021learning].
  5. 5.
    Search Space Reduction - Another line of work aims to directly reduce the search space by eliminating decisions unlikely to appear in an optimal solution. For example, in the travelling salesman problem, ML models can predict irrelevant arcs for pruning. Studies following this approach include [fischetti2019machine, fitzpatrick2021learning, grassia2019learning, lauri2019fine, larocca2024combining, li2025learning, spieckermann2025reduce, sun2021generalization, tahir2021improved, tayebi2022learning, vaclavik2018accelerating].
  6. 6.
    Constraint Programming - Constraint Programming (CP) solves combinatorial problems through systematic constraint propagation and search. Several works leverage supervised learning to guide CP solvers by predicting constraint satisfaction or search decisions, including [selsam2018learning, sun2024enhancing, xu2018towards].
  7. 7.
    Algorithm Configuration - ML is also used to decide whether specific algorithmic components should be applied to a given instance. Examples include predicting when to apply Dantzig–Wolfe decomposition [kruber2017learning], when to use scaling techniques [berthold2021learning], whether to apply cutting planes beyond the root node [berthold2206learning], and whether to linearise the objective in mixed-integer quadratic programmes [bonami2018learning]. In all cases, algorithm runs are used to generate labelled training data.
Problems in this category may be either single-stage or sequential. Single-stage settings include tasks such as predicting high-quality initial solutions for warm-starting. In contrast, sequential settings require a series of interdependent decisions, such as branching choices along a B&B tree. [Nair2020SolvingMI] addresses both perspectives. In addition, some approaches rely on a single prediction to inform multiple downstream decisions. For example, [khalil2022mip] predicts likely variable assignments at the root node of a binary linear program and reuses this information for subsequent node selection.
3.3.2 End-to-End Learning
- End-to-end learning refers to a modelling paradigm that tightly integrates ML and optimisation within a single framework. In this setting, the learning model either directly predicts solutions to optimisation problems, corresponding to the AIT, or produces partially specified optimisation models whose parameters are inferred from raw or structured input data, corresponding to the PPT.
Parameter Prediction Task
- Early work in this direction includes [donti2017task] and [elmachtoub2022smart]. In particular, [elmachtoub2022smart] introduce the Smart “Predict, then Optimise” (SPO) framework, which trains ML models to predict uncertain coefficients in linear objective functions. As the original SPO loss is non-convex and difficult to optimise, the authors propose a convex surrogate, SPO+, which enables efficient training. Subsequent studies apply the SPO and SPO+ losses across a range of applications, including last-mile delivery, energy systems, and ship management [chu2023data, demirovic2020dynamic, sang2022electricity, tian2023smart].
[donti2017task] extend regret minimisation to stochastic convex optimisation problems and rely on implicit differentiation of the regret objective. More recently, [baty2024combinatorial, greif2024combinatorial, jungel2025learning, rautenstrauss2025optimization] propose learning pipelines enriched with optimisation layers and apply them to dynamic vehicle routing problems. These works typically rely on full-information offline solutions as supervision and aim to minimise regret with respect to these benchmarks. Notably, [greif2024combinatorial] and [rautenstrauss2025optimization] employ DAgger-based data collection strategies to improve robustness.
Related ideas appear in [parmentier2021learning], which studies a single-machine scheduling problem using structured prediction [nowozin2011structured] to approximate hard instances via easier surrogate problems whose parameters are learned. [sinclair2023hindsight] leverage hindsight or full-information solutions to estimate expert values and propose a hindsight learning algorithm, demonstrating its effectiveness across multiple problem domains.
Action Imitation Task
- An early attempt to directly predict solutions to difficult combinatorial problems, such as the Travelling Salesman Problem (TSP), is due to [vinyals2015pointer], who propose pointer networks trained in a supervised manner on solved instances. Their approach demonstrates limited generalisation to instances of varying sizes. Building on this idea, [li2018combinatorial] and [joshi2019efficient] develop supervised models for static CO problems.
[kotary2024learning] propose a supervised learning framework for predicting optimal solutions based on observable features correlated with unknown parameters of a parametric optimisation problem. [chen2023two] address the Flexible Job Shop Scheduling Problem using a two-stage supervised framework that imitates a constraint programming solver, first predicting machine assignments and then job start times, followed by post-hoc feasibility restoration.
Several works extend action imitation to dynamic or stochastic settings. [larsen2018predicting] study the Load Planning Problem by imitating second-stage decisions of a two-stage optimisation model using full-information supervision. [ozarik2024machine] learn driver routing behaviour in last-mile delivery from historical data. [qi2023practical] consider dynamic inventory management under demand uncertainty, training supervised models on offline full-information solutions. [pham2023prediction] apply regression-based learning to dynamic radiotherapy appointment scheduling. More recently, [gawas2024ildagger] study the Employee Notification Timing Problem and use DAgger to train policies that imitate a range of offline expert models.
ML-augmented optimisation and end-to-end learning thus represent two dominant paradigms for combining ML with optimisation. While ML-augmented approaches embed learning within specific algorithmic components such as branching, cutting, or search space reduction, end-to-end methods aim to predict decisions or optimisation parameters directly through unified learning pipelines. Across both paradigms, supervision plays a central role, yet the nature of the expert varies substantially, ranging from full-information solutions and heuristics to intermediate solver states. These choices have significant implications for learning performance, generalisability, and computational cost. Understanding how experts are defined and utilised is, therefore, crucial for algorithm design. We next introduce a taxonomy of expert types to clarify these distinctions and to guide the development of IL methods for SDPs.
4 Expert Taxonomy
In IL, an expert is a policy, or more generally a decision-making function, that specifies the desired behaviour in a sequential decision problem. The expert guides the learning agent by providing action demonstrations for given states, typically assuming access to an optimal or near-optimal strategy under a known or implicit objective. Related terms used in the literature include oracle, demonstrator, optimal policy, ground truth, reference policy, and target. We broadly distinguish between two categories of experts.
  • Human Expert – Human experts rely on experience and contextual understanding to make decisions under uncertainty. In dynamic settings such as vehicle dispatching or healthcare scheduling, they implicitly incorporate real-time information, operational constraints, and institutional knowledge. For example, [ozarik2024machine] study last-mile delivery by imitating driver decisions that capture practical insights not explicitly represented in optimisation models. While human decisions can be nuanced, they may also be inconsistent or subject to cognitive bias. Moreover, collecting large and reliable datasets from human experts is often difficult due to limited availability and high cost.
  • Algorithmic Expert – Algorithmic experts generate decisions using mathematical models, optimisation solvers, or rule-based and heuristic methods. Such experts can typically produce large volumes of consistent demonstrations. Their design depends on how the decision objective is specified and how uncertainty, denoted by $\boldsymbol{\xi}$, is handled. The vector $\boldsymbol{\xi}$ captures exogenous factors such as demand, travel times, or arrival processes that are not fully known at the time of decision making.
Given the complexity of real-world SDPs and the need for timely decisions, model-based algorithmic experts are often more suitable for generating demonstrations in practice.
Ideally, for an SDP, the optimal action at decision epoch $k$ is obtained by solving (7) where the expectation is taken with respect to future uncertainty.
$a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\mathbb{E}_{\xi}\Big\{C_{k}(x_{k},a_{k})+V_{k+1}(x_{k+1})\,\big|\,x_{k}\Big\},$ (7)
For many real-world business applications, computing such optimal actions, even offline, is often impractical due to problem complexity. Consequently, it is common in operations research to design algorithms that produce near-optimal solutions—decisions that are not globally optimal but are sufficiently close to be useful in practice. When such solutions can be computed within reasonable time limits, they can serve as expert demonstrations for IL.
Despite the widespread use of approximate or surrogate experts, the literature lacks a clear and systematic taxonomy for defining experts and their roles within IL. To address this gap, we propose a structured taxonomy of expert types to guide the design and application of IL methods in SDPs.
4.1 Expert Decision Models under Uncertainty
We further classify algorithmic experts according to how they handle uncertainty $\boldsymbol{\xi}$ at the time of decision making. The treatment of $\boldsymbol{\xi}$ determines whether an expert behaves myopically or anticipates future outcomes, and whether it is relatively simple or highly adaptive. We outline five classes of algorithmic experts, distinguished by how uncertainty is incorporated into their decision models.
  1. 1.
    Myopic Expert - A myopic expert selects actions based solely on immediate costs, effectively ignoring future uncertainty. It treats $\boldsymbol{\xi}$ as unavailable or irrelevant and bases decisions only on the current state. This leads to simple models and fast solution times, but often results in poor long-term performance due to the lack of anticipation.
  2. 2.
    Deterministic Expert - The deterministic expert replaces uncertain parameters $\boldsymbol{\xi}$ with fixed point estimates $\hat{\boldsymbol{\xi}}$, such as mean or nominal values, and solves the resulting deterministic optimisation problem. While straightforward to implement, this approach may perform poorly when actual realisations deviate substantially from the assumed estimates.
  3. 3.
    Full-Information (Hindsight) Expert - The full-information expert assumes perfect foresight and solves the optimisation problem using the realised values $\tilde{\boldsymbol{\xi}}$ of all uncertain parameters. Although infeasible in real-world operations, this expert serves as a useful benchmark, as it yields the best possible decision in hindsight for each instance.
  4. 4.
    Two-Stage Stochastic Expert - In the two-stage stochastic framework, uncertainty $\boldsymbol{\xi}$ is not observed at the initial decision stage. The expert first selects a here-and-now (first stage) action and subsequently adapts through recourse decisions (second stage) once uncertainty is revealed. The cost function can be divided into two parts, where the first part $Q_{1}(a)$ represents the first stage cost and $Q_{2}(a,\xi)$ denotes the second stage cost. This structure captures limited anticipation by committing to an initial action before uncertainty is known and then optimally adjusting with recourse actions after the uncertainty is revealed. Thus, it produces more robust supervision than previous experts. It remains computationally manageable while preserving expressive power, which is why they are widely adopted in stochastic programming.
  5. 5.
    Multi-Stage Stochastic Expert - Multi-stage stochastic experts operate in fully dynamic settings where uncertainty unfolds gradually over time. Decisions are updated at each stage based on newly revealed information, allowing policies to adapt as $\boldsymbol{\xi}$ evolves. While this formulation closely reflects real-world decision making under uncertainty, it is computationally expensive to solve exactly, particularly when uncertainty is high-dimensional or difficult to model.
Table 2 summarises the objective structure and treatment of uncertainty for each expert type. Here, $C$ denotes the total cost function and $\mathcal{A}$ represents the set of all feasible actions.
Table 2: Summary of Expert Models
Policy/Action Model Type Objective Uncertainty
$\pi^{*m},a^{*m}$ Myopic $\operatorname*{arg\,min}\limits_{a\in\mathcal{A}}\{C(a)\}$ No uncertainty
[…]
$\pi^{*t},a^{*t}$ Two-Stage Stochastic $\operatorname*{arg\,min}\limits_{a\in\mathcal{A}^{1}}\{C(a,\xi)=Q_{1}(a)+\mathbb{E}_{\xi}[Q_{2}(a,\xi)]\}$ Uncertainty Set
$\pi^{*},a^{*}$ Multi-Stage Stochastic $\operatorname*{arg\,min}\limits_{a\in\mathcal{A}}\mathbb{E}_{\xi}\{C(a,\xi)\}$ Full Uncertainty
Most of the literature in SDP relies on either myopic or full-information experts for imitation, depending on the problem structure and availability of future information. In settings where uncertainty is absent—such as static or fully observable optimisation problems—a deterministic expert is typically employed. Since all relevant parameters are known with certainty, the expert can compute a single optimal solution using standard optimisation techniques.
4.2 Expert Optimality
Even when employing the expert types defined above, the underlying optimisation problem may remain difficult to solve. Many deterministic optimisation problems are $\mathcal{NP}$-hard, making exact solution approaches computationally impractical in many real-world settings. As a result, practitioners often rely on heuristics or impose stopping criteria when using optimisation solvers. We therefore classify experts according to the optimality guarantees of the models they solve.
  • Task-Optimal Expert – A task-optimal expert solves its underlying optimisation model to optimality, assuming the problem is tractable and can be solved exactly within reasonable computational limits. Such experts are most applicable when efficient solution methods are available. For example, a strong branching expert evaluates linear programming relaxations for all fractional variables and selects the branching variable that yields the greatest objective improvement [Nair2020SolvingMI, alvarez2016online]. Optimal solutions may be obtained using specialised solvers, including mixed-integer programming solvers [sun2022learning], SAT solvers [sun2024enhancing], and dedicated solvers such as Concorde for the TSP [sun2021generalization]. We denote the policy of a task-optimal expert by $\pi^{*\cdot}$ and the corresponding optimal action by $a^{*\cdot}$, where $\cdot$ refers to the expert type in Table˜2.
  • Approximate Expert – An approximate expert produces actions that trade off solution quality for computational efficiency, aiming to generate near-optimal decisions. Common approximation strategies include:
    • Time limits: terminating the solver after a fixed time budget and returning the best solution found [li2025learning, baty2024combinatorial, pham2023prediction, sun2021generalization].
    • Optimality gaps: stopping once a predefined optimality gap is achieved [larsen2018predicting].
    • Objective-based refinement: applying alternative formulations or local improvements to progressively improve solution quality [ding2020accelerating].
    • Heuristics: directly using high-quality heuristic methods, such as hybrid genetic search for the prize-collecting TSP [baty2024combinatorial], or evaluating a finite set of candidate actions and selecting the best-performing one [song2020general].
    We denote the policy of an approximate expert by $\pi^{\cdot}$ and the corresponding action by $a^{\cdot}$.
4.3 Expert Interaction
Depending on the cost, availability, and reliability of the expert, different strategies can be used to incorporate expert supervision into IL.
  • Learning without interaction - In this non-iterative setting, the expert is queried once to label a batch of observed states, after which the learning model is trained offline. This approach is appropriate when expert queries are expensive or time-consuming, as it minimises interaction. Most studies in the literature adopt this strategy by generating a fixed dataset of expert-labelled states. However, as discussed in Section˜3, one-shot learning is vulnerable to poor generalisation when the learned policy encounters states that were not present in the expert dataset. We also refer to this setting as Isolated learning.
  • Learning with interaction - Interactive expert querying is commonly used to enrich training data and mitigate distribution shift by exposing the learner to a broader range of states. [he2014learning] employ the DAgger algorithm to iteratively query an expert when learning node selection policies for branch-and-bound. Similarly, [song2020general] propose a DAgger-based approach to learn search policies for both $A^{*}$ and branch-and-bound algorithms in integer programming. Several recent works adopt iterative expert interaction in combinatorial optimisation settings, including [gawas2024ildagger, greif2024combinatorial, Nair2020SolvingMI, rautenstrauss2025optimization].
    While effective, the standard DAgger algorithm can be computationally expensive due to frequent expert queries. To reduce this cost, alternative interaction schemes have been proposed [labassi2022learning, rautenstrauss2025optimization]. Related work also considers online learning settings, where models are updated incrementally as new data are collected, which likewise requires expert interaction. Examples include [alvarez2016online, khalil2016learning, gasse2019exact, lu-2020].
Figure˜3 summarises the literature on expert models, highlighting distinctions between task-optimal and approximate experts, as well as between one-shot and interactive interaction strategies. The figure indicates that stochastic expert models and interactive learning approaches remain relatively uncommon, likely due to the additional methodological and computational complexity they introduce.
Figure 3: Classification of literature.
LiteratureMyopicDeterministicHindsight2-Stage Stochastic[alvarez2017machine], [furian2021machine], [gupta2022lookback],[khalil2017learning], [kraul2023machine], [morabit2021machine],[Nair2020SolvingMI], [ojha2023optimization], [paulus2022learning],[shen2022enhancing], [sonnerat2021learning], [sun2022learning][gasse2019exact], [khalil2016learning], [vaclavik2018accelerating][guaje2024machine], [huang2023searching], [lin2022learning],[paul2025data], [pereira2022learning], [zarpellon2021parameterizing][lu-2020][demirovic2020dynamic], [elmachtoub2022smart], [julien2024machine],[jungel2025learning], [kotary2024learning], [niroumandrad2024learning],[qi2023practical], [sinclair2023hindsight], [wang2023imitation][he2014learning], [gawas2024ildagger], [labassi2022learning],[rautenstrauss2025optimization][baty2024combinatorial], [cai2024multi], [chu2023data],[donti2017task], [hottung2020deep], [larsen2018predicting],[li2025learning], [pham2023prediction], [sang2022electricity],[tian2023smart], [yilmaz2021study][song2020general][detassis2021teaching], [fitzpatrick2021learning], [joshi2019efficient],[kaempfer2018learning], [lauri2019fine], [li2018combinatorial],[liu2022learning], [lodi2020learning], [morabit2022machine],[parmentier2021learning], [selsam2018learning], [spieckermann2025reduce],[sun2021generalization][gawas2024ildagger] [chen2023two], [ding2020accelerating], [gerbaux2025machine],[grassia2019learning], [han2023gnn], [hizir2025large],[khalil2022mip], [larocca2024combining], [matsuoka2019machine],[mossina2019multi], [tayebi2022learning], [vinyals2015pointer],[xavier2021learning], [xu2018towards], [yaakoubi2020flight],[yang2023learning][greif2024combinatorial],[Nair2020SolvingMI][abbasi2020predicting][gawas2024ildagger]OptimalApproximateOne-ShotInteractive
4.4 Labelling strategy
Most existing studies rely on an expert to provide a single label for each state. However, for certain types of experts, it is possible to obtain multiple labels for the same state within one or more interactions.
  • Single label per state - This is the standard approach, in which the expert is queried once to obtain a single target action for a given state.
  • Multiple labels per state - Expert actions may vary due to approximation methods or inherent stochasticity in the underlying models. To improve robustness, multiple labels can be collected for the same state and combined using aggregation techniques such as majority voting or averaging. This process reduces variability and mitigates the impact of expert error. Multiple labels can be obtained using two mechanisms.
    1. 1.
      Single call - A single interaction with the expert yields multiple candidate actions for a given state. This occurs, for example, when solving an integer program using a mathematical programming solver that returns multiple feasible solutions. In [khalil2022mip], CPLEX is allowed to run for 60 minutes to generate a solution pool for each instance, terminating after identifying 1,000 solutions with objective values within 10% of the best solution found.
    2. 2.
      Multiple calls - Alternatively, repeated queries to the expert may be used to collect several actions for the same state. This approach is common when using deterministic experts in dynamic stochastic programming, where the optimal action depends on sampled future scenarios [gawas2024ildagger, greif2024combinatorial].
4.5 Number of Experts Used
Algorithms can also be categorised based on the number of experts involved and how their expertise is integrated during learning. We identify two primary categories:
  • Single Expert- The conventional setting in IL, where a single, typically optimal expert provides guidance. The learner assumes the expert is both reliable and inexpensive to query, consistently supplying accurate actions for all states. This simplifies learning but may limit adaptability in complex or diverse environments. Most literature uses a single expert to build the dataset.
  • Multi-Expert- In some cases, multiple experts with complementary specialisations are available. Their actions can be aggregated into a composite "meta-expert", combining diverse expertise to enhance robustness and flexibility across different aspects of the task. [song2020general] uses two different experts to get demonstrations. First, they use an expensive expert to gather initial expert demonstrations, which are used to train initial IL policies. Next, a simpler retrospective expert is used to augment the dataset. [rautenstrauss2025optimization] uses a set of non-optimal policies to label states for which the true hindsight expert is expensive. These states generally represent initial decision epochs in the horizon. [sun2023mega] introduces MEGA-DAgger, a variant of DAgger designed for interactive learning with multiple imperfect experts.
4.6 Trajectory Control
When experts are employed within an iterative learning framework, they can also be used to control trajectories and correct errors made by a previously learned model. This approach ensures that diverse states are seen that are likely to be encountered during deployment. The following control strategies are commonly used:
  • Total Expert Control - The expert governs all states visited within a learning trajectory. A typical example is pure supervised learning without iteration, where the expert labels all states once without further interaction.
  • Diminishing Expert Control - The learner initially relies heavily on the expert, but gradually reduces this dependence as its performance improves. This approach is exemplified by the basic DAgger algorithm, also referred to as Vanilla DAgger [ross2011reduction]. This type of control is most used in interactive experts in [gawas2024ildagger], [greif2024combinatorial], [song2020general].
  • Conditional Expert Control - The expert is invoked selectively based on performance criteria, such as the confidence of the learned model action falling below a threshold, as in SAFE DAgger [zhang2016query]. Alternatively, an external policy may guide trajectories when the true expert is expensive or imperfect.
Moreover, to reduce reliance on costly expert calls, the expert is queried only for critical cases, such as novel or poorly handled states. This selective querying improves sample efficiency and focuses learning on high-impact scenarios. [menda2019ensembledagger] and [hoque2021lazydagger] describe the Ensemble DAgger and Lazy DAgger algorithms, where the expert is selectively queried based on the discrepancy measure between the expert and learned policy actions. However, such variants of the DAgger algorithms have not been experimented with for SDP in operations research.
Using the proposed expert taxonomy, we introduce and evaluate an Adaptive DAgger-based algorithm that maps each dimension of the taxonomy to an explicit design choice in the learning procedure. This framework enables a structured comparison of expert configurations within an action imitation (AIT) learning setting.
5 Algorithm
In this section, we present a general-purpose adaptive algorithm for learning from experts in sequential decision problems. The proposed algorithm, shown in Algorithm˜3, explicitly incorporates the expert taxonomy introduced in Section˜4. A summary of the algorithm parameters is provided in Table˜3.
Table 3: List of parameters for Algorithm˜3
Parameter Description
$\mathcal{D}$ Dataset containing state–expert action pairs
$H$ Number of times to simulate current learned policy $\pi_{i}$
$\mathcal{E}=\{\pi^{*}_{1},\pi^{*}_{2},\dots\}$ Set of available experts
$K$ Episode trajectory length
$F_{j}$ Function that determines the number of expert calls at a given state
$J$ Number of expert calls in a given state
$F_{e}$ Function that selects which expert to query
$\kappa$ Stopping criterion for the expert
AA Aggregation function for combining multiple expert actions
DRDR Decision rule for selecting the executed action
STOP() Stopping criterion
The algorithm is based on the DAgger framework, which follows an iterative policy learning paradigm inspired by online learning and assumes access to one or more experts. Let $\mathcal{E}$ denote the set of available experts. At each learning iteration, indexed by $i$, the learner retrains its policy using all states encountered so far, together with the corresponding expert-labelled actions. This iterative process allows the learner to correct earlier mistakes and progressively improve their performance.
The procedure starts from an initial policy $\hat{\pi}_{0}$, which may be obtained from expert demonstrations or derived from a heuristic. This policy is then executed for $H$ simulated episodes using the decision rule, each consisting of $K$ decision epochs. At each epoch, the learner observes the current state and may query an expert up to $J$ times, as determined by the function $F_{j}$. Each query returns an expert action, potentially from different experts in $\mathcal{E}$, decided by the expert-selection function $F_{e}$.
We denote an expert action by $a^{\cdot}_{ijk}$, where $i$ indexes the learning iteration, $j$ the expert call within the epoch, and $k$ the decision epoch. For clarity, we omit the episode index in this notation. This may be task-optimal or approximate depending on the complexity of the underlying optimisation problem. The parameter $\kappa$ specifies the list of expert stopping criteria, thereby controlling the quality of the returned action. The resulting state–action pairs are stored in the dataset before the transition to the next state occurs.
After completing $H$ episodes, the dataset is augmented with the newly collected expert-labelled samples, and a new policy $\hat{\pi}_{i+1}$ is trained. This process is repeated until a stopping condition is satisfied, yielding progressively refined policies. The following paragraphs relate the components of the expert taxonomy to Algorithm˜3.
1
2
3Function DAgger():
4   Initialize dataset $\mathcal{D}\leftarrow\emptyset$; initial policy $\hat{\pi}_{0}$; iteration counter $i\leftarrow 0$;
5   learning routine $\textsc{Train}(\cdot)$;
6 while STOP() do
    // Simulate $H$ episodes using current policy $\pi_{i}$
7
8    for $h\in\{1,\dots,H\}$ do
       // Sample a $K$-step trajectory
9       for $k\in\{1,\dots,K\}$ do
10            Observe state $x_{k}$;
          // Make $J$ expert calls
11          $J\leftarrow F_{j}(x_{k})$;
12          for $j\in\{1,\dots,J\}$ do
13             $\pi^{*}\leftarrow F_{e}(x_{k},\mathcal{E},j)$;
14             ${a}^{\cdot}_{ijk}\leftarrow\pi^{*}(x_{k},\kappa)$;
15
16            end for
17
18          $\hat{a}_{ik}\leftarrow A\big(\{{a}^{\cdot}_{ijk}\}_{j=1}^{J}\big)$;
19          $\mathcal{D}\leftarrow\mathcal{D}\cup\{(x_{k},\hat{a}_{ik})\}$;
20
21          $a_{k}\leftarrow DR(x_{k},\hat{a}_{ik},\hat{\pi}_{i})$;
22
23          $x_{k+1}\leftarrow X^{M}(x_{k},a_{k},\xi_{k})$;
24
25         end for
26
27      end for
28
29    $i\leftarrow i+1$;
30    $\hat{\pi}_{i+1}\leftarrow\textsc{Train}(\mathcal{D})$;
31
32   end while
33
34 end
Algorithm 3 Adaptive DAgger Algorithm
We characterize the imitation learning framework with respect to the taxonomy: the expert type is given by the set $\mathcal{E}$, which may include myopic, deterministic, full-information, or two-stage stochastic models; expert optimality, governed by a stopping criterion $\kappa$ that controls solution accuracy and computational effort; and expert interaction, determined by the number of iterations $I$, yielding a sequence of policies $\{\hat{\pi}_{1},\dots,\hat{\pi}_{I}\}$. Additional design choices include the labelling strategy $F_{j}$, which specifies when and how often expert queries are issued and how multiple actions $a^{*}_{ijk}$ are aggregated via $A$; the number of experts $|\mathcal{E}|$, with expert selection handled by $F_{e}$; and trajectory control, defined by a decision rule $DR$.
6 Application to Generalised Assignment Problem
This section presents the problem setting and experimental framework used to evaluate the proposed imitation learning methodology. We first introduce the dynamic physician-to-patient assignment problem, which serves as a representative sequential decision problem under uncertainty and allows for systematic instantiation of the expert taxonomy. We then describe the experimental design, including expert configurations, learning setups, and baseline policies, and report a comprehensive empirical evaluation of the proposed framework. The experiments are designed to assess how different expert types, levels of optimality, and interaction strategies influence learning performance, data efficiency, and computational cost.
6.1 Generalised Assignment Problem
The classical Generalised Assignment Problem (GAP) has been extensively studied since the seminal work of [balachandran1976integer]. The problem consists of assigning a set of items (or jobs) to a set of agents (or machines) such that each job is assigned to exactly one agent, agent capacity constraints are respected, and the total assignment cost or profit is optimised. GAP has found applications in a wide range of domains, including job scheduling [balachandran1976integer], routing [fischetti2019machine], and patient admission scheduling [ceschia2011local]. A standard integer programming formulation of the GAP is given by (8)-(11). Here, $c_{ij}\in\mathbb{R}_{+}$ denotes the cost incurred by assigning job $j$ to agent $i$, $d_{ij}\in\mathbb{R}_{+}$ is the resource consumption of agent $i$ when processing job $j$, $L_{i}\in\mathbb{R}_{+}$ is the capacity of agent $i$, and $x_{ij}$ is a binary decision variable equal to $1$ if job $j$ is assigned to agent $i$ and 0 otherwise.
$\min\quad$ $\sum_{i=1}^{m}\sum_{j=1}^{n}c_{ij}x_{ij}$ (8)
s.t. $\sum_{i=1}^{m}x_{ij}=1,\quad$ $\forall j=1,\dots,n,$ (9)
$\sum_{j=1}^{n}d_{ij}x_{ij}\leq L_{i},\quad$ $\forall i=1,\dots,m,$ (10)
$x_{ij}\in\{0,1\},\quad$ $\forall i=1,\dots,m,\ \forall j=1,\dots,n.$ (11)
6.1.1 Physician-to-Patient Assignment
We consider a dynamic physician-to-patient assignment (PPA) problem as a specific application of the GAP. In a clinical setting, appointment requests arrive sequentially over a booking horizon until a fixed cut-off time. The clinic is staffed by a set of physicians providing services during a given session. Let $\mathcal{P}=\{1,\dots,P\}$ denote the set of physicians. Each physician $p\in\mathcal{P}$ has a regular session duration $T$ (in minutes) and a maximum appointment capacity $L_{p}$.
Appointment requests arrive independently over time and are handled individually by a receptionist. At each arrival, the receptionist must decide whether to accept or reject the request based on the current state of the system. If the request is accepted, the receptionist must assign the patient to a physician. Each patient $k$ has a preferred physician $p_{k}$ among a set of eligible physicians $\mathcal{P}_{k}\subseteq\mathcal{P}$. Each patient is characterised by a service duration $d_{k}\in\mathbb{R}_{+}$, which is estimated at the time of the call, and a priority class $r_{k}\in\{1,2\}$. The arrival process is such that calls from priority 1 patients arrive near the end of the horizon.
Rejecting a patient of priority $r_{k}$ incurs a penalty $c^{\text{rej}}_{r_{k}}$, with $c^{\text{rej}}_{1}>c^{\text{rej}}_{2}$. In addition, assigning a patient to an undesired but eligible physician incurs a preference penalty $c^{\text{pref}}_{r_{k}}$. The primary source of uncertainty in the system arises from the random arrival times and sequences of patients within a session. This creates a trade-off between accepting lower-priority patients early and preserving capacity for potential high-priority patients who may arrive later. Furthermore, assigning patients to undesired physicians may reduce future capacity, potentially forcing rejections later in the booking horizon. The resulting problem is an SDP in which each decision epoch corresponds to a patient arrival, and the action is to assign the patient to an eligible physician or to reject the request. The offline mixed-integer programming formulation, denoted by $(\textbf{{MIP}}_{\textbf{PPA}})$, is provided in the appendix.
We adopt the PPA problem to demonstrate the expert taxonomy. It requires anticipating future arrivals and strategically reserving capacity for high-priority patients, a core challenge in sequential decision making under uncertainty. This structure allows us to distinguish between myopic, deterministic, and stochastic experts based on how they incorporate future information. At the same time, the underlying CO formulations remain relatively tractable, ensuring that the computational bottleneck of repeated expert queries in DAgger is manageable. Feasibility is also straightforward to enforce through capacity constraints, enabling clear evaluation of learned policies without extensive post-processing. Finally, PPA requires rapid decision-making at each patient call-in, making it well-suited for assessing the benefits of learned policies that replace expensive optimisation with fast deployment-time decisions. We emphasise that our goal is not to propose a new solution method for the PPA itself, but rather to use it as a representative SDP for systematically analysing expert design choices within IL.
6.1.2 MDP Formulation
We now formulate the PPA problem as an MDP.
Decision Epochs – Let $\mathcal{K}=\{0,1,\dots,K\}$ denote the set of decision epochs, where each epoch corresponds to the arrival of a patient call-in. Here, $K$ denotes the total number of requests within a session.
State Variable – At decision epoch $k$, the system state is given by $x_{k}=\big(x^{k}_{\text{patient}},x^{k}_{\text{physician}},x^{k}_{\text{eligibility}}\big)$. The component $x^{k}_{\text{patient}}=[d_{k},r_{k}]$ represents the characteristics of the current patient. The component $x^{k}_{\text{physician}}=[x^{k}_{p}]_{p\in\mathcal{P}}$ captures the status of all physicians, where $x^{k}_{p}=\big({1}_{p=p_{k}},x^{k}_{pl},x^{k}_{p1},x^{k}_{pw}\big).$ Here, ${1}_{p=p_{k}}$ indicates whether physician $p$ is the patient preferred physician, $x^{k}_{l}$ denotes the remaining assignment capacity, $x^{k}_{p1}$ is the number of priority-1 patients already assigned, and $x^{k}_{pw}$ is the current workload. A list of indicators gives the eligibility vector functions $x^{k}_{\text{eligibility}}=[{1}_{p\in\mathcal{P}_{k}}]_{p\in\mathcal{P}}.$
Decision Variable – At each decision epoch $k$, the receptionist selects an action $a_{k}\in\mathcal{A}_{k}=\{0\}\cup\mathcal{P}_{k}$, where $a_{k}=0$ corresponds to rejecting the request and $a_{k}=p$ corresponds to assigning the patient to physician $p$. A physician $p$ is eligible only if $d_{k}+x^{k}_{pw}\leq L_{p}$, ensuring feasibility of all actions in $\mathcal{A}_{k}$.
Exogenous Information – The exogenous information at epoch $k$ corresponds to the characteristics of the next arriving patient and is denoted by $\xi_{k}=[d_{k+1},r_{k+1},p_{k+1},\mathcal{P}_{k+1}].$
Transition Function – The transition function $X^{M}(\cdot)$ maps the system from state $x_{k}$ to $x_{k+1}$ according to $x_{k+1}=X^{M}(x_{k},a_{k},\xi_{k}).$ Specifically, for the selected physician $p=a_{k}$,
  • $x^{k+1}_{pl}=x^{k}_{l}-{1}$,
  • $x^{k+1}_{p1}=x^{k}_{p1}+{1}_{r_{k}=1}$,
  • $x^{k+1}_{pw}=x^{k}_{pw}+d_{k}$.
Cost Function – At each decision epoch, a cost $C_{k}(x_{k},a_{k})$ is incurred, defined as
$C_{k}(x_{k},a_{k})=\begin{cases}c^{\text{rej}}_{r_{k}},&\text{if }a_{k}=0,\\[6.0pt] c^{\text{pref}}_{r_{k}},&\text{if }a_{k}\neq 0\text{ and }a_{k}\neq p_{k},\\[6.0pt] 0,&\text{otherwise.}\end{cases}$
6.2 Results
To demonstrate how different expert properties influence learning outcomes, we evaluate the performance of Algorithm˜3 on the physician-to-patient assignment (PPA) problem. All expert optimisation models are solved using Gurobi. The full experimental codebase is publicly available at GitHub. All experiments are conducted on a machine equipped with an Intel Xeon Gold 6258R CPU running at 2.70  GHz. The remainder of this section is organised as follows. Section˜6.2.1 describes the experimental setup, including instance generation, evaluation metrics, baseline policies, and the computational environment. Section˜6.2.2 presents and discusses the experimental results in detail.
6.2.1 Experimental Setup
This subsection describes the instance generation procedure and the configuration of the learning models and baseline methods.
Data
- We begin by describing the parameters used to generate instances of the PPA. The proposed approach is evaluated on synthetically generated instances constructed under controlled conditions. The total number of patient calls arriving at the clinic is denoted by $K$, as each patient corresponds to a decision epoch. $K$ is sampled from a normal distribution with mean $\mu=100$ and standard deviation $\sigma=8$. Each patient is assigned a priority class: class 1 (high priority) with probability $0.3$, and class 2 (regular priority) otherwise.
To capture the tendency of high-priority (emergency) patients to call later in the day, we generate an arrival score for each patient from a beta distribution, $\mathrm{Beta}(a,b)$. Specifically, arrival scores are drawn with parameters $(a=3,b=1)$ for class 1 patients and $(a=1,b=1)$ for class 2 patients. Patients are then ordered according to these arrival scores, ensuring that high-priority patients are more likely to appear towards the end of the scheduling horizon.
Appointment durations are assumed to be known at the time of the call and are sampled from a lognormal distribution. For class 1 patients, durations are generated using parameters $\mu=3$ and $\sigma=0.8$, whereas for class 2 patients the parameters are $\mu=2.3$ and $\sigma=0.3$. We consider a setting with $P=4$ physicians. To construct each patient’s eligibility set, physician weights are first sampled from a Dirichlet distribution with parameter vector $[0.4,0.3,0.15,0.15]$. The top $k$ physicians with the largest weights are then selected, where $k$ is drawn uniformly from $\mathcal{U}\{1,\dots,P\}$. We describe this process in detail in the appendix.
Rejection penalties are set to $c^{\text{rej}}_{1}=200$ for priority 1 patients and $c^{\text{rej}}_{2}=50$ for priority 2 patients. Assigning a patient $i$ to a non-preferred but eligible physician incurs a penalty of $c^{\text{pref}}_{r_{i}}=0.1\times c^{\text{rej}}_{r_{i}}$.
ML Models
- We construct a simulation model of the PPA using the data generation parameters described above. At each decision epoch, the learned policy is invoked to select an action. For each decision epoch, we generate a training example of the form $(x_{k},\hat{a}_{k})$, where $x_{k}$ denotes the concatenation of the state features and $\hat{a}_{k}$ is the target action provided by the expert at epoch $k$. The learning model maps the current state features to a decision that either assigns the patient to an eligible physician or rejects the request.
As the learning architecture, we employ a multilayer perceptron (MLP) consisting of an input layer, a single hidden layer with 64 neurons, and an output layer. Normalised state features are provided as input, and the MLP outputs a probability distribution over the available actions, namely assignment to each eligible physician or rejection. The final action is selected based on this predicted probability vector.
Baselines
- We compare the learned policies against three baseline methods. The first is a greedy policy, denoted by $\pi_{\text{greedy}}$, which assigns an arriving patient to their preferred physician whenever possible. If the preferred physician is unavailable, the patient is randomly assigned to any eligible physician. This policy also serves as a myopic expert.
The second baseline, denoted by $\pi^{t}$, corresponds to a two-stage stochastic expert. At each decision epoch, this policy samples 30 future demand scenarios. These scenarios are generated by simulating future arrivals conditioned on the current system state. The expert solver is then invoked to compute an action by solving the resulting two-stage stochastic optimisation problem over the generated scenarios.
The final baseline is the exact optimal full-information policy, denoted by $\pi^{*f}$, computed separately for each test instance. This policy represents an upper bound on achievable performance and serves as a benchmark for assessing the quality of learned and heuristic policies.
6.2.2 Performance Evaluation
We evaluate multiple configurations of expert models using Algorithm˜3 and compare their performance against the baseline policies. All methods are tested on a dataset of 1,000 instances. For each policy, we report the average objective value, the total number of rejected patients by priority class, and the total number of assignments to non-preferred physicians, aggregated across both priority classes. Each expert model is given a time limit of 30 seconds to compute an action at each decision epoch.
Baselines
- Table˜4 summarises the performance of the baseline policies. The table reports the average cost alongwith the % of rejections for each priority and % of undesirable assignments by each policy. The main difference between the performance of all policies is due to priority 1 patient rejections. The greedy policy $\pi_{\text{greedy}}$ performs poorly, primarily due to a high percentage of rejections among priority 1 patients. In contrast, the two-stage stochastic policy $\pi^{t}$ achieves substantially better performance by reducing rejections across both priority 1 patients. This improvement comes at the cost of a modest increase in assignments to non-preferred physicians.
It is important to note that $\pi^{t}$ is computationally expensive despite the time limit. With 30 sampled future scenarios, the policy requires approximately 25 hours to complete 10 episodes when executed in parallel. $\pi^{*f}$ shows the best achievable performance with significantly low priority 1 patient rejections.
Table 4: Average performance metrics for $\pi_{greedy}$, $\pi^{t}$ and $\pi^{*f}$
Policy Cost Priority 1 Rejected % Priority 2 Rejected % Undesirable Assignments (%)
$\pi_{greedy}$ 3047.48 39.7 20.2 22.3
$\pi^{t}$ 1785.58 10.4 26.4 27.5
$\pi^{*f}$ 1271.25 1.4 26.2 25.5
Learning without interaction (Isolated Learning)
- We first report results obtained without expert interaction. The hyperparameter settings used for Algorithm˜3 are summarised in Table˜5 and correspond to the base configuration of the algorithm. Any deviations from this configuration are explicitly stated in the relevant experimental subsections. Setting $I=1$ implies that only a single outer iteration is performed, resulting in the training of a single learned policy.
Table 5: Hyper Parameters for Algorithm˜3
Hyperparameters
STOP() $I=1$
$F_{j}(.)$ 1
$\kappa$ 30 seconds
Decision Rule Vanilla DAgger
$\lambda$ 1
$\beta$ 0.8
$F_{e}(\cdot)$ $\pi^{*f}$, $\pi^{*d}$ $\pi^{t}$
We consider three expert policies as targets: the full-information expert ($\pi^{*f}$), the deterministic expert ($\pi^{*d}$), and the two-stage stochastic expert ($\pi^{*t}$). We cannot include the multistage expert $\pi^{*}$, as it is too difficult to obtain even for small instances of the problem. Each expert is queried once per state. The corresponding learned policies are denoted by $\hat{\pi}^{d}$, $\hat{\pi}^{*d}$, $\hat{\pi}^{*d}$. To construct the dataset of expert demonstrations, we simulate a total of $H=4000$ episodes. Using Algorithm˜3, we collect approximately 400,000 expert demonstrations for each of the three expert targets. Due to the imposed 30-second time limit on expert optimisation, only the two-stage stochastic expert $\pi^{*t}$ terminates early $(\textbf{{MIP}}_{\textbf{SPPA}})$ and may return a non-optimal action. As a result, the learned policy is ultimately trained using demonstrations generated by $\pi^{t}$. In contrast, both the full-information and deterministic experts can solve their respective optimisation models ($\textbf{{MIP}}_{\textbf{PPA}}$) to optimality within the allotted time. We also note that employing the full-information expert $\pi^{*f}$ within Algorithm˜3 is equivalent to solving deterministic instances of the PPA using a mixed-integer programming solver in one shot.
To employ the two-stage stochastic expert ($\pi^{t}$), it is necessary to specify the number of future scenarios generated at each decision epoch. In this study, we use a fixed number of scenarios across all epochs. Alternative designs could allow the number of scenarios to vary over time—for example, by generating more scenarios in later stages, where the two-stage problem may be easier to solve, and decisions are likely to be more critical with respect to the utilisation of remaining capacity. We evaluate three configurations using 5, 15, and 30 scenarios. Finally, data collection is parallelised using 10 threads, with 10 episodes simulated simultaneously. All training times and results reported in this study are based on this parallel execution setup.
We first analyse the effect of the number of simulated episodes on the performance of the learned policy, to understand how the size of the demonstration dataset influences learning quality. The results are shown in Figure˜4. For the full-information expert, increasing the number of demonstrations does not lead to a clear or consistent improvement in performance. In contrast, policies trained using demonstrations from the deterministic expert exhibit a steady performance improvement as the number of demonstrations increases.
A plausible explanation is that the deterministic expert encounters a wider range of system states, as it does not select actions that are optimal with respect to the realised trajectory. Consequently, the learned model is exposed to a more diverse state distribution, including states that would not typically be visited under an optimal policy. This increased diversity improves the generalisation capability of the learned policy when deployed. The black line in Figure˜4 represents the performance of the baseline policy $\pi^{t}$. However, learning from a two-stage stochastic expert with multiple scenarios substantially outperforms learning from a full-information and deterministic expert based on a single scenario.
Figure 4: Test performance of learned models with different numbers of episodes NN
Table˜6 reports detailed results for the best-performing policies learned using all three experts. The higher costs of $\hat{\pi}^{*f}$, $\hat{\pi}^{*d}$ relative to $\pi^{t}$ baseline are primarily attributable to an increased number of rejected priority 1 patients under the policy learned from the full-information expert. The table also reports the total training time required to simulate 4,000 episodes. It is important to note that the full-information expert can label all states within an episode in a single optimisation run by solving the $\textbf{{MIP}}_{\textbf{PPA}}$ once, and therefore does not require the use of Algorithm˜3 for data generation. In contrast, the deterministic expert must generate and solve a new optimisation instance at every state within an episode, which substantially increases the overall data collection time.
The learned policies using ${\pi}^{t}$ as the target markedly reduce patient rejections. Among the configurations considered, training with 30 scenarios yields the best performance and comes closest to the baseline policy $\pi^{t}$, as expected. These improvements come at the cost of increased training time, since multiple scenarios must be generated and the corresponding multi-scenario optimisation problem $(\textbf{{MIP}}_{\textbf{SPPA}})$ must be solved. The results indicate substantial performance improvements as the size of the training dataset increases. Performance gains are expected to plateau once the number of episodes exceeds approximately 2,000.
Table 6: Average performance metrics for the best $\hat{\pi}^{\cdot}$
Learnt Policy Scenarios Cost
Priority 1
Rejected %
Priority 2
Rejected%
Undesirable
Assignments (%)
Train Time (hours)
Full-Information ($\hat{\pi}^{f}$) 1 2031.34 13.2 29.5 25.8 0.5
Deterministic ($\hat{\pi}^{d}$) 1 1975.72 11.8 29.8 25.5 3.25
Two-Stage ($\hat{\pi}^{t}$) 5 1890.35 12.1 26.4 27.5 8.65
15 1822.45 10.8 26.6 27.4 32.47
30 1801.85 10.4 26.8 27.2 65.94
Table˜7 reports detailed statistics on optimality gaps and solution times under the imposed 30-second time limit, aggregated over approximately 400,000 expert calls for the two-stage stochastic expert using $\textbf{{MIP}}_{\textbf{SPPA}}$. Most instances are solved to optimality or near optimality, as indicated by the 99% values. As the number of scenarios increases, a modest rise in the average optimality gap is observed. When gaps are non-zero, this is typically due to the solver identifying the optimal solution but requiring additional time to certify optimality. Overall, these results indicate that the stochastic formulation remains tractable and provides high-quality expert demonstrations.
Average solution times increase with the number of scenarios. As a result, deploying $\pi^{t}$ with 30 scenarios for real-time decision making would require approximately 5.09 seconds per decision on average, whereas the learned policy produces decisions within milliseconds. Solution times also vary over the planning horizon. As shown in Figure˜5, early decision epochs exhibit larger optimality gaps and longer solve times, reflecting the larger problem size at the beginning of an episode. Under a 30-second time limit, solution times at early epochs can approach 20 seconds on average, further underscoring the value of fast and reliable learned policies. On the other hand, solving the $\textbf{{MIP}}_{\textbf{PPA}}$ for the full-information and deterministic expert requires, on average, 0.2 seconds per instance.
Table 7: Statistics across expert ($\pi^{t}$) calls
Scenarios Average Gap (%) 99% Gap (%) Average Solve Time (secs) 99% Solve Time (secs)
5 0 0 0.66 6.41
15 0.004 0.16 2.51 30
30 0.013 0.29 5.09 30
Figure 5: Average Gap and Average solve time across decision epoch
Learning with interaction
- We next allow interaction with the expert in order to iteratively learn decision policies. This requires modifying the stopping criterion of the learning procedure. Specifically, we impose a stopping criterion of $I=400$ learning iterations. In each iteration $i$, the current learned policy $\pi_{i}$ is executed for $H=10$ episodes. This results in a total of 4,000 simulated episodes used to construct the training datasets, matching the total number of episodes used in the isolated learning setting.
Since the single-scenario experts solve their underlying optimisation problems to optimality relatively quickly, they are able to support a large number of learning iterations. Upon completion of the learning process, we simulate all learned policies and select the policy that achieves the lowest average cost on the test instances.
Figure˜6 reports the average cost across learning iterations for all expert types. The horizontal axis corresponds to the learned policy $\pi_{i}$ obtained after iteration $i$. The results show a consistent decrease in average cost as additional data are incorporated through expert interaction. Performance metrics for the best learned policies are summarised in Table˜8. We also show the average objectives of the best models obtained through isolated learning, indicated by dotted lines in the figure.
The models learned through interaction achieve lower average costs compared with those obtained via isolated learning, as reported in Table˜6. Table˜8 also reports the percentage improvement relative to the isolated learning models. In addition, policies learned with interaction reject fewer patients overall. No significant performance difference is observed between policies learned from the full-information expert and those learned from the deterministic expert.
As shown in Figure˜6, learning performance begins to saturate after approximately 200 iterations, although the best-performing models are obtained at later iterations. We would like to point out that the best models obtained through isolated learning required 4000 training episodes. Using interaction, the learned model improves on those values within 2000 episodes. This highlights the effect of interaction. Compared with the full-information and deterministic expert settings, a stronger and more consistent decreasing trend in total cost is observed as learning progresses. Among all configurations, the policy trained using 30 scenarios achieves the best performance. It yields a marginal improvement over the policy learned without interaction, while requiring a significantly smaller number of expert demonstrations, coming very close to the baseline policy $\pi^{t}$.
Training times are also reported in Table˜8. Under interactive learning, the training time associated with the deterministic expert more than doubles. For the full-information expert, training time increases substantially due to the use of the vanilla DAgger decision rule. Under this rule, the states encountered during learning differ from those observed by a pure full-information expert, which necessitates solving the $\textbf{{MIP}}_{\textbf{PPA}}$ at each newly encountered state. Similarly, the two-stage stochastic expert also sees a considerable increase in training times.
Figure˜6(b) shows the same plot with respect to training time in minutes. Within the first 6 hours of training, we see that the Two-Stage Stochastic Expert with 5 scenarios shows better test performance than those with 15 and 30. Similarly, the one with 15 scenarios is better than 30 scenarios for the next 10 hours. This suggests adding more states and learning more efforts in the short term is better. But eventually, if one allows more time, the one with higher scenarios gives better models.
Figure 6: Test performance of different learned models
(a) With respect to learned models
(b) With respect to time
[…]
Table 8: Average performance metrics for the best model with interaction
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected%
Priority 2
Rejected%
Undesirable
Assignments (%)
Train Time (hours)
Full-Information ($\hat{\pi}^{f}$) 1 261 1951.76 3.9 10.9 29.8 26.9 16.40
Deterministic ($\hat{\pi}^{d}$) 1 371 1943.85 1.6 11.3 29.3 25.6 14.77
Two-stage Stochastic ($\hat{\pi}^{t}$) 5 389 1844.75 2.4 10.65 27.4 27.2 23.19
15 301 1804.59 0.9 10.1 27.2 27.0 59.22
30 251 1795.93 0.3 10.0 27.1 27.2 101.83
Comprising Optimality of Experts
Figure˜6 shows that the two-stage stochastic expert achieves the best performance, albeit at the cost of substantial training times. To reduce the overall computational burden, we investigate whether training time can be lowered by relaxing the optimality requirement of the two-stage stochastic expert. As illustrated in Figure˜5, the first few decision epochs are the most computationally demanding and account for the majority of total training time. In comparison, a fixed 30-second time limit appears overly conservative for later epochs, where problem instances are smaller and easier to solve.
To address this imbalance, we introduce an additional stopping criterion based on the MIP gap. Specifically, we set the tolerance to $2\%$, meaning that the expert terminates once this optimality gap is reached. As shown in Figure˜5, the observed gaps are, on average, already well below this threshold. We therefore set $\kappa=2\%$ and repeat the same experiment using the two-stage stochastic expert under this relaxed optimality condition.
Table˜9 reports the best policies obtained across all three scenario settings using the two-stage stochastic expert. The learned policies match the performance of those in Table˜8, which were trained under a fixed 30-second time limit. Notably, the policy trained with 30 scenarios achieves even stronger results than its counterpart in Table˜8. One possible explanation is that solving the two-stage problem to near optimality as suggested by the MIP gaps obtained, may bias the expert targets toward the sampled scenarios. In contrast, terminating based on an MIP gap criterion may yield targets that better approximate the underlying multistage stochastic optimisation problem represented by $\pi^{*}$. We also observe a substantial reduction in training times—particularly for the 15- and 30-scenario settings—without any deterioration in policy performance.
Table 9: Average performance metrics for the best model with interaction with MIP gap 0.2%
Expert Scenarios Best Model Cost
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments (%)
Train Time (hours)
Two-stage Stochastic ($\hat{\pi}^{t}$) 5 381 1845.1 11.11 26.98 27.14 20.95
15 376 1811.2 10.55 26.68 27.44 23.01
30 370 1786.0 9.8 27.15 27.2 26.56
Figure˜8 further illustrates the average MIP gaps and the time required by the expert to compute target actions at each decision epoch. The plateau observed in the gap curve indicates the decision epochs for which the approximation criterion $\kappa$ becomes binding. This effect is mirrored by a noticeably flatter profile in the average solve time across all three scenario settings. As a result, the overall training time is substantially reduced.
Figure 7: Average Gap and Average solve time at each epoch with MIP gap 0.2% across scenarios
Figure 8: Test performance of learnt models for different MIP gap settings
Additionally, we varied the MIP gap parameter when training exclusively with the two-stage stochastic expert using 30 scenarios. Specifically, we considered four values: $0.01$, $0.02$, $0.05$, and $0.1$. The corresponding test performance is shown in Figure˜8. We observe similar learning behaviour for gap limits of $0.01$, $0.02$, and $0.05$, whereas the $0.1$ setting exhibits a noticeable decline in performance, suggesting that the resulting demonstrations are of lower quality. The best-performing policies under each configuration are summarised in Table˜10. These results indicate that, although the two-stage stochastic problem can be computationally demanding, effective policies can still be learned by moderately relaxing optimality. Additional experimental results exploring variations of other algorithmic parameters are provided in the appendix.
Table 10: Average performance metrics for the best model with interaction with 30 scenarios
Expert MIP Gap Best Model Cost
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments (%)
Train Time (hours)
Two-stage Stochastic ($\hat{\pi}^{t}$) 0.01 377 1783.5 9.93 26.94 27.15 33.41
0.02 370 1786.0 9.8 27.15 27.2 26.56
0.05 341 1788.2 9.43 27.82 27.11 24.97
0.1 397 1810.7 8.82 29.3 24.85 24.84
7 Open Research Directions
Despite rapid progress in imitation learning for CO under uncertainty, several important research directions remain open.
  1. 1.
    Adaptive Expert Query Strategies — Although DAgger-style interaction improves robustness by mitigating compounding errors, expert queries remain computationally expensive in many optimisation settings. An important open direction is the development of principled mechanisms that decide both when to query and which expert to consult, based on criteria such as model uncertainty, state novelty, or expected performance gain. Integrating selective sampling and confidence-aware querying [zhang2016query, menda2019ensembledagger] from classical IL literature into stochastic optimisation frameworks is particularly promising.
  2. 2.
    Multi-Expert Supervision — The taxonomy shows that experts offer complementary strengths: deterministic experts provide efficiency, whereas stochastic experts deliver stronger supervision. As illustrated in Algorithm˜3, heterogeneous experts can be combined adaptively based on the observed state. Learning meta-policies for state-dependent expert selection could reduce computational burden while maintaining solution quality. In the PPA, early decision epochs are especially demanding for two-stage stochastic models; employing a milder expert at these stages can yield substantial time savings. A similar idea appears in ambulance routing [rautenstrauss2025optimization].
  3. 3.
    Learning under Structured Expert Imperfection — Most studies assume bounded optimality gaps or heuristic approximations but offer limited theoretical insight into how expert suboptimality propagates through imitation learning. Establishing performance guarantees that relate expert approximation quality to learned policy performance remains an important open problem, particularly in stochastic settings.
  4. 4.
    Sequential Prediction Architectures — Decision spaces are large in CO problems, making accurate end-to-end prediction difficult. A promising approach is to decompose decisions into sequential stages, with separate models predicting different components. For example, in the Dynamic Inventory Routing Problem [greif2024combinatorial], one may first select customers and then determine their visiting order. Training coordinated models, each guided by a specialised expert, may improve scalability and predictive performance.
8 Conclusion
This study examined the role of the expert in imitation learning for SDPs under uncertainty. Using the PPA case study, we compared expert configurations in terms of interaction, optimality, and computational effort. Policies trained with the two-stage stochastic expert using 30 scenarios achieved the strongest performance, both with and without interaction, while interaction further improved policy quality. Deterministic and full-information experts enable rapid generation of demonstrations but tend to produce lower-quality supervision, whereas two-stage stochastic experts provide stronger guidance at higher computational cost.
For the PPA, the two-stage formulation remains tractable with a moderate number of scenarios, but computational effort increases substantially in early decision epochs. Relaxing optimality requirements via an MIP gap limit significantly reduced training time without degrading learned policy performance, highlighting the trade-off between expert quality and scalability.
From a computational perspective, learned ML policies amortise the cost of expert computation, producing fast and stable decisions at deployment. Heuristic multi-scenario optimisation methods can reduce computation but may be biased towards sampled scenarios and perform poorly under rare events. In contrast, learned policies aggregate information across instances, improving robustness and efficiency. The Aggregated Deterministic Expert, which solves scenarios independently and aggregates decisions [gawas2024ildagger, greif2024combinatorial], offers a practical compromise: while not matching two-stage stochastic performance, it consistently outperforms single-scenario deterministic and full-information experts.
Overall, the findings emphasise the central role of expert design in imitation learning for sequential decision making. By formalising expert characteristics through a structured taxonomy and embedding them within a DAgger-based framework, this work provides a unified perspective for interpreting prior results and guiding expert selection. Future work may further explore adaptive expert-query strategies, richer stochastic expert models, and extensions to larger-scale optimisation problems.
Appendix
Appendix
Appendix A Background
A.1 MDP
Given the core components from Section˜2, we further describe the value function and policy in detail here.
  • Value function - The value function represents the cumulative cost over the planning horizon. For a cost-minimisation objective, the optimal value function from an initial state $x_{0}$ is given by
    $V^{*}(x_{0})=\min_{a}\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},a_{k})\,\big|\,x_{0}\Bigg\},$ (12)
    We omit the discount factor for simplicity.
  • Policy ($\pi$) - A policy $\pi$ is a mapping from states to admissible actions $a_{k}=\pi(x_{k})$. Policies may be represented in tabular form or as parameterised functions, for example, with parameter $\theta$. The value of a policy $\pi$ is defined as
    $V^{\pi}(x_{0})$ $=\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},\pi(x_{k}))\,\big|\,x_{0}\Bigg\}$ (13)
    The optimal policy $\pi^{*}$ is then defined as
    $\pi^{*}$ $=\operatorname*{arg\,min}_{\pi}\mathbb{E}\Bigg\{\sum_{k=0}^{K}C_{k}(x_{k},\pi(x_{k}))\,\big|\,x_{0}\Bigg\}$ (14)
    $=\operatorname*{arg\,min}_{\pi}V^{\pi}(x_{0}).$ (15)
    In the MDP literature, $V^{\pi^{*}}(x_{k})$ is referred to as the optimal cost-to-go from state $x_{k}$ at epoch $k$.
Given a value function, the optimal policy can be recovered via one-step greedy minimisation. Value functions for a fixed policy can be computed using dynamic programming methods such as value iteration or policy iteration [puterman2014markov]. While these approaches yield exact solutions, they are often computationally infeasible for large-scale problems due to the curse of dimensionality [powell2007approximate].
In some applications, interest lies only in the optimal action at a specific epoch $k$. The optimal action is given by
$a_{k}^{*}=\operatorname*{arg\,min}_{a_{k}}\Big\{C_{k}(x_{k},a_{k})+\mathbb{E}[V^{*}_{k+1}(x_{k+1})]\,\big|\,x_{k}\Big\},$ (16)
where $V^{*}_{k+1}(x_{k+1})$ denotes the optimal value of the subsequent state.
Appendix B MIP formulations
We first provide an MIP formulation for PPA for the static problem as well as the stochastic problem with $\Omega$ scenarios.
B.1 Deterministic MIP formulation
We give an MIP formulation for the PPA below. We denote the patients by $k$ and the physicians by $p$.
Table 11: Model, Parameters, and Variables
Sets
$\mathcal{P}$ set of all physicians
$\mathcal{K}$ set of all patients
$\mathcal{P}_{k}$ set of all eligible physicians for patient $k$
$\{1,2\}$ priority classes
Parameters
$P$ number of physicians
$K$ number of patients
$t_{k}$ duration of patient $k$
$r_{k}$ priority of patient $k$
$p_{k}$ preferred physician of patient $k$
$T$ total work capacity for physician $p$
$L_{p}$ total number of appointments allowed for physician $p$
$c_{r_{k}}^{\text{rej}}$ rejection cost for patient $k$
$c_{r_{k}}^{\text{pref}}$ penalty for assigning an undesired physician to patient $k$
Variables
$a_{kp}\in\{1,0\}$ $k\in\mathcal{K}$, $p\in\mathcal{P}$ 1 if patient $k$ is assigned to physician $p$
$u_{k}\in\{1,0\}$ $k\in\mathcal{K}$ 1 if patient $k$ is rejected
The objective function (17) is the sum of the total rejection cost and the penalty for all non-preferred assignments. Constraint (18) ensures that all patients are either assigned to a physician or rejected. Constraints (19) and (20) enforce capacity limits on the number of appointments and total workload, respectively.
$\textbf{{MIP}}_{\textbf{PPA}}:=\min\quad\underbrace{\sum_{k\in\mathcal{K}}c^{\text{rej}}_{r_{k}}u_{k}}_{\text{Rejection cost}}\;+\;\underbrace{\sum_{k\in\mathcal{K}}\sum_{p\in\mathcal{P}}c^{\text{pref}}_{r_{k}}a_{kp}}_{\text{Non-preferred assignment penalty}}$ (17) s.t. $\sum_{p\in\mathcal{P}_{k}}a_{kp}+u_{k}=1$ $\forall k\in\mathcal{K}$ (18) $\sum_{k\in\mathcal{K}}a_{kp}\leq L_{p}$ $\forall p\in\mathcal{P}$ (19) $\sum_{k\in\mathcal{K}:p\in\mathcal{P}_{k}}t_{k}\,a_{kp}\leq T$ $\forall p\in\mathcal{P}$ (20) $a_{kp}\in\{0,1\}$ $\forall k\in\mathcal{K},\ \forall p\in\mathcal{P}$ (21) $u_{k}\in\{0,1\}$ $\forall k\in\mathcal{K}$ (22)
B.2 Two-stage Stochastic MIP formulation
Table˜12 introduces additional notation for the two-stage formulation, where variables and parameters are augmented with a scenario index $\omega$. All other variables and sets retain the same meaning as in Table˜11. We use $\omega$ to denote a scenario. In a two-stage formulation, the first patient is common across all scenarios and is denoted by $k=1$.
Table 12: Model, Parameters, and Variables
Sets
$\Omega$ set of all scenarios
$\mathcal{K}_{\omega}$ set of all patients in scenario $\omega$
$\mathcal{P}_{k\omega}$ set of all eligible physicians for patient $k$ in scenario $\omega$
Parameters
$K_{\omega}$ number of patients in scenario $\omega$
$t_{k\omega}$ duration of patient $k$ in scenario $\omega$
$r_{k\omega}$ priority of patient $k$ in scenario $\omega$
$p_{k\omega}$ preferred physician of patient $k$ in scenario $\omega$
$c_{r_{k\omega}}^{\text{rej}}$ rejection cost for patient $k$ in scenario $\omega$
$c_{r_{k\omega}}^{\text{pref}}$ penalty for assigning an undesired physician to patient $k$ in scenario $\omega$
Variables
$a_{kp\omega}\in\{1,0\}$ 1 if patient $k$ is assigned to physician $p$ in scenario $\omega$
$u_{k\omega}\in\{1,0\}$ 1 if patient $k$ is rejected in scenario $\omega$
$\textbf{{MIP}}_{\textbf{SPPA}}:=\min\ \frac{1}{|\Omega|}\left[\sum_{\omega\in\Omega}\sum_{k\in\mathcal{K}_{\omega}}c^{\text{rej}}_{r_{k\omega}}u_{k\omega}+\sum_{\omega\in\Omega}\sum_{k\in\mathcal{K}_{\omega}}\sum_{p\in\mathcal{P}_{k\omega}}c^{\text{pref}}_{r_{k\omega}}a_{kp\omega}\right]$ (23) s.t. $\sum_{p\in\mathcal{P}_{k\omega}}a_{kp\omega}+u_{k\omega}=1$ $\forall\omega\in\Omega,\ \forall k\in\mathcal{K}_{\omega}$ (24) $\sum_{k\in\mathcal{K}_{\omega}:p\in\mathcal{P}_{k\omega}}a_{kp\omega}\leq L_{p}$ $\forall\omega\in\Omega,\ \forall p\in\mathcal{P}$ (25) $\sum_{k\in\mathcal{K}_{\omega}:p\in\mathcal{P}_{k\omega}}t_{k\omega}\,a_{kp\omega}\leq T$ $\forall\omega\in\Omega,\ \forall p\in\mathcal{P}$ (26) $a_{1p\omega}=a_{1p\omega^{\prime}}$ $\forall\omega\neq\omega^{\prime}$ (27) $u_{1\omega}=u_{1\omega^{\prime}}$ $\forall\omega\neq\omega^{\prime}$ (28) $a_{kp\omega}\in\{0,1\}$ $\forall\omega\in\Omega,\ \forall k\in\mathcal{K}_{\omega},\ \forall p\in\mathcal{P}$ (29) $u_{k\omega}\in\{0,1\}$ $\forall\omega\in\Omega,\ \forall k\in\mathcal{K}_{\omega}$ (30)
The objective (23) and constraints (24)–(26) mirror those of $\textbf{{MIP}}_{\textbf{PPA}}$. Constraints (27) and (28) enforce non-anticipativity of the first-stage decisions across scenarios.
Appendix C Dirichlet-based eligibility construction
Eligible physicians are generated using a patient-level preference model based on the Dirichlet distribution. Let $\mathcal{P}=\{1,\dots,P\}$ denote the physician set. For each arriving patient, a probability vector
$\mathbf{w}=(w_{1},\dots,w_{P}),\qquad w_{p}\geq 0,\ \sum_{p\in\mathcal{P}}w_{p}=1,$
is specified as a global baseline over physicians. In the experiments with $P=4$, we use $\mathbf{w}=[0.4,0.3,0.15,0.15]$. A concentration parameter $\alpha>0$ controls how strongly patient preferences concentrate around $\mathbf{w}$. We form the Dirichlet parameter vector
$\boldsymbol{\alpha}=\alpha\,\mathbf{w},$
and sample a patient-specific taste vector
$\mathbf{z}=(z_{1},\dots,z_{P})\sim\mathrm{Dirichlet}(\boldsymbol{\alpha}).$
Each component $z_{p}$ represents the relative affinity of the current patient for physician $p$, with $\sum_{p}z_{p}=1$. Larger values of $\alpha$ produce tastes that are closer to the baseline $\mathbf{w}$, whereas smaller values yield more dispersed and heterogeneous tastes across patients. In our implementation, $\alpha=25$ is used when the baseline weights are enabled, and $\alpha=1$ is used when all physicians are treated uniformly (that is, $w_{p}=1/P$ for all $p$).
Given $\mathbf{z}$, the eligible set size is drawn as
$K\sim\mathcal{U}\{k_{\min},\dots,k_{\max}\},$
and the eligible physician set for the patient is defined by selecting the $K$ physicians with the largest taste values:
$\mathcal{P}_{k}=\operatorname{TopK}(\mathbf{z},K).$
Equivalently, we construct an eligibility indicator $E_{p}\in\{0,1\}$ for each physician $p$, where $E_{p}=1$ if and only if $p\in\mathcal{P}_{k}$. A safety check enforces that at least one physician is eligible, which is guaranteed by construction for $K\geq 1$.
After constructing $\mathcal{P}_{k}$, the preferred physician is selected from within the eligible set using the taste vector restricted to $\mathcal{P}_{k}$. Specifically, the preferred physician is sampled with probabilities proportional to $\{z_{p}:p\in\mathcal{P}_{k}\}$, so that physicians with higher taste are more likely to be preferred while maintaining feasibility with respect to eligibility.
Appendix D Results
We present some results on additional experiments.
D.1 Aggregated Deterministic Expert
The aggregated deterministic expert [gawas2024ildagger], [greif2024combinatorial] is motivated by settings in which solving a full multi-scenario stochastic optimisation problem is computationally expensive or unstable under tight time limits. Instead of jointly optimising across scenarios, this approach solves multiple deterministic instances independently—each corresponding to a sampled future scenario—and then aggregates the resulting actions. This decomposition reduces solver complexity, enables parallelisation, and provides robustness by smoothing over scenario-specific noise. Aggregated deterministic experts therefore offer a practical compromise between purely deterministic and fully stochastic experts, especially when high-quality stochastic solutions are difficult to obtain within the available computational budget.
D.1.1 Learning without interaction
For a fixed number of scenarios, the deterministic expert, with $F_{e}(\cdot)$ set to $\pi^{*f}$, can be used to solve each scenario instance independently, after which the resulting solutions are aggregated to produce a single target action. Instance generation follows the same procedure as for the two-stage stochastic expert. Under this strategy, aggregation effectively serves as a heuristic approximation of a two-stage stochastic problem. Relative to the base configuration of Algorithm˜3, the primary modification is that the number of expert actions returned by $F_{j}(\cdot)$ is equal to the number of scenarios considered.
Figure˜9 illustrates the impact of the number of simulated episodes on the performance of the learned policy. Performance improves as the number of episodes increases during the initial phase of training. However, beyond approximately 500 episodes, additional data provide limited benefit, and in some cases, a marginal deterioration in performance is observed.
Figure 9: Aggregated Deterministic Expert
Table˜13 reports results across all scenario configurations considered. For aggregation, we convert the optimal actions into frequency vectors and use these as training targets for the ML model. This yields a smoother, continuous target representation, as opposed to the binary action targets used with the other expert configurations. Across all scenario counts, performance in terms of total cost remains broadly similar, indicating that increasing the number of scenarios offers limited additional benefit. Compared with policies learned from the two-stage stochastic expert, the aggregated deterministic policies perform worse overall.
Moreover, training time is substantially higher for the aggregated deterministic expert. This is because the expert must solve multiple optimisation problems sequentially to generate each target action, rather than solving a single multi-scenario formulation. As a result, the total training time exceeds that of the two-stage stochastic expert by a significant margin.
Table 13: Average performance metrics for Aggregated Deterministic Expert
Learnt Policy Scenarios Average Cost
Priority 1
Rejected%
Priority 2
Rejected%
Undesirable
Assignments%
Train Time (hours)
Deterministic ($\pi^{*d}$) 5 1944.87 11.8 29.0 25.4 15.90
15 1950.09 11.9 29.0 25.5 50.63
30 1958.12 12.0 29.0 25.4 97.18
D.1.2 Learning with interaction
- Figure˜10 illustrates the performance of learned policies obtained using different numbers of aggregated scenarios, while Table˜14 reports the performance of the best-performing policies. No significant performance differences are observed across the scenario counts considered. Notably, the best-performing policy is obtained using aggregation over only five scenarios. This observation is consistent with the spread across scenario configurations shown in Figure˜9. Furthermore, the learned models perform marginally better than those obtained through isolated learning, while requiring a smaller number of expert demonstrations.
Figure 10: Aggregated Deterministic Expert
Overall, the learned policies marginally outperform those trained using the full-information and deterministic experts. However, they do not achieve the same level of performance as policies learned from the two-stage stochastic expert. In addition, increasing the number of aggregated scenarios to 15 or 30 does not yield meaningful performance improvements relative to the isolated learning results reported in Table˜13.
Table 14: Average performance metrics for the best model using the Aggregated Deterministic Expert with interaction
Expert Scenarios Best Model Cost % Improvement
Priority 1
Rejected
Priority 2
Rejected
Undesirable
Assignments
Train Time(hours)
Aggregated Deterministic 5 366 1938.69 0.3 11.2 29.5 25.3 29.16
15 359 1941.93 0.4 11.5 29.3 25.5 64.42
30 190 1939.165 0.9 11.4 29.3 25.3 116.45
D.1.3 Inference
Overall aggregated deterministic expert provides an alternative to the two-stage stochastic expert, but for the PPA, it does not yield superior policies. Moreover, aggregation requires repeated expert calls, significantly increasing training time. Empirical results show that, for the same number of scenarios, Algorithm˜3 with an aggregated deterministic expert incurs a higher computational cost than the two-stage stochastic expert. That said, as demonstrated in [gawas2024ildagger], aggregated deterministic experts can outperform two-stage stochastic experts in settings where the latter are difficult to solve to high quality within practical time limits.